### v2

# Attractors for the Wave Equation with Nonlinear Energy Critical Sources and Dissipation

Abstract
The wave equation with energy-critical nonlinear sources and dissipation in a bounded 3 D domain is considered. It is shown that the resulting dynamical system admits a global attractor. Under the additional assumption of strong monotonicity of the damping at the origin, the said attractor is also finite dimensional and smooth. In order to handle “energy criticality” of both sources and the damping, the methods used depend on enhanced dissipation [7], energy identity for weak solutions [21], an adaptation of Ball’s method [1], and the theory of quasi-stable systems [13].
Keywords: Wave equation; global attractor; nonlinear damping; nonlinear source; critical exponents.
MSC: 35B33; 35B40; 35B41; 35L05; 35L70.
1 Introduction
Let $\Omega\subset\mathbb{R}^{3}$, where $\Omega$ is a bounded domain with a sufficiently smooth boundary $\Gamma=\partial\Omega$. The following classical energy critical wave model is under consideration in this work:
$\left\{\begin{aligned} &u_{tt}-\Delta u+g(u_{t})+f(u)=h,\quad\mbox{in}\quad\Omega\times\mathbb{R}^{+},\\ &u|_{\Gamma}=0,\quad\mbox{on}\quad\Gamma=\partial\Omega,\\ &u(x,0)=u_{0}(x),\quad u_{t}(x,0)=u_{1}(x),\quad x\in\Omega,\end{aligned}\right$. (1.1)
where $g$ is a nonlinear monotone damping of critical, quintic nature: $|g(s)|\leq M|s|^{5}$ and the source $f$ is also a nonlinear critical source with quintic upper bound: $|f(s)|\leq M|s|^{5},\,|s|>1$.
A study of wave equation in 3 dimensions with a quintic source have attracted a lot of activities. Quintic term is critical for the wave equation in 3 dimensions, thus a host of mathematical issues at the level of well-posedness of weak solutions. In fact, while local well-posedness can be obtained via several methods, global solvability has been a source of major challenges. And here is the reason why. The strategy used for establishing existence of weak solutions to undamped wave equation relies heavily on Strichartz estimates which allow to trade differentiability into suitable integrability. However, Strichartz estimates are constructed for each solutions [so called Strichartz solution] $u$ on time interval $[0,T]$. These estimates loose uniformity with respect to bounded set of initial data when time $T\rightarrow\infty$, unless nonlinearity is subquintic. The above predicament, particularly sensitive to long time behavior of the models, stimulated enormous amount of research. A good account is presented in [20]. To address this issue, methods relaying on “nonconcentration” of singularity in S-estimates have been developed [9]. These allow to extend local solutions to global ones. Additional arguments relying on the so called backward regularity [29] lead to S-solutions which form a dynamical system. These methods also rely on finite speed of propagation which can be also proved in the presence of linearly bounded dissipation-[20]. However, in the case of highly nonlinear dissipation, proving globality and uniqueness of solutions by using the above described method appears challenging.
This brings us to the topic of the present paper. We consider damping also of critical exponent with an enhanced dissipation. The problem goes back to [27] and [6, 7, 5] where such dynamics has been considered. In fact, the idea was to use the enhanced dissipation in order to establish well-posedness, particularly uniqueness of weak solutions corresponding to critical and supercritical sources.. The estimates in these references provide an initial road map to our analysis, which eventually culminates with the theory of coherent attracting sets. Thus, our goal is to take the problem to the next level by considering long time behavior-in particular existence and a structure of attracting sets. This will be accomplished by developing appropriate tools for establishing asymptotic smoothness [particularly in the degenerate case when $g^{\prime}(0)=0$], finite-dimensionality and smoothness of the attractors. The challenge here is due to double criticality of the source and dissipation. Criticality of the source makes potential energy functional $\int_{\Omega}F(u(x))dx$ with $F^{\prime}=f$ noncompact, thus preventing methods of compensated compactness from applicability. Criticality of the dissipation implies $g(u_{t})\in H^{-1}(\Omega)$, but $g(u_{t})\notin H^{-\alpha}(\Omega),\alpha<1$. The latter is the critical regularity for variational solutions to be well defined with test functions compactly embedded in $H^{1}(\Omega)$. Our final result establishes an existence of global attractor which, moreover, is finite-dimensional and smooth.
2 Functional Setting, Assumptions and Main Results
2.1 Functional Setting and the Assumptions Imposed
First of all, we will define some notations that will be used throughout the work. We denote by $H^{s}(\Omega)$ the $L^{2}$- based Sobolev’s space of the order $s$ and by $H^{s}_{0}(\Omega)$ the closure of $C^{\infty}_{0}(\Omega)$ in $H^{s}(\Omega)$. In particular, the inner products and norms in the spaces $L^{2}(\Omega)$ and $H^{1}_{0}(\Omega)$ will be denoted, respectively, by
$\left(u,v\right)=\int_{\Omega}uvdx,\quad\|u\|^{2}=\int_{\Omega}|u(x)|^{2}dx,$
$((u,v))=\left(\nabla u,\nabla v\right),\quad\|u\|^{2}_{H^{1}_{0}(\Omega)}=\|\nabla u\|^{2}.$
We consider the Laplacian operator $\Delta$ defined by the triplet $\{H^{1}_{0}(\Omega),L^{2}(\Omega),a(u,v)\}$ where $a(u,v)$ is the bilinear continuous form on $H^{1}_{0}(\Omega)\times H^{1}_{0}(\Omega)$ given by
$a(u,v)=\left(\nabla u,\nabla v\right),\quad u,v\in H^{1}_{0}(\Omega).$
It is known from Spectral Theorem for self-adjoint operators that the domain of $\Delta$ is given by $D[(-\Delta)]=H^{2}(\Omega)\cap H^{1}_{0}(\Omega)$ and there exists a complete orthonormal family of $L^{2}(\Omega)$, $\{\omega_{j}\}_{j\in\mathbb{N}}$, made of eigenfunctions of $\Delta$ under Dirichlet boundary conditions.
$\Delta\omega_{j}=\lambda_{j}\omega_{j},\;j=1,\cdots,\quad\mbox{and}\quad 0<\lambda_{1}\leq\lambda_{2}\leq\cdots,\lambda_{j}\rightarrow+\infty\quad\mbox{as}\quad j\rightarrow+\infty.$
We can also define the powers $(-\Delta)^{s}$ of $\Delta$, for $s\in\mathbb{R}$. The spaces
$H^{2s}=D[(-\Delta)^{s}],s\in\mathbb{R}\quad\text{ (with Dirichlet boundary conditions)}$
are Hilbert spaces with the norms $\|u\|_{H^{2s}}=\|(-\Delta)^{s}u\|$. We have $H^{2s_{1}}\subset H^{2s_{2}}$, $\forall s_{1},s_{2}\in\mathbb{R},$ $s_{1}\geq s_{2}$, where each space is dense in the following one and the injection is compact if $s_{1}>s_{2}$. Moreover
$\|(-\Delta)^{s_{2}}u\|\leq\lambda_{1}^{s_{2}-s_{1}}\|(-\Delta)^{s_{1}}u\|\quad\mbox{if}\quad s_{1}\geq s_{2}.$
We identify $D[(-\Delta)^{0}]=L^{2}(\Omega)$ with its dual, which leads to the identification $H^{-2s}=D[(-\Delta)^{-s}]$ with the dual space of $H^{2s}=D[(-\Delta)^{s}]$, $s>0$. In particular, we have $H^{0}=L^{2}(\Omega)$, $H^{1}=H^{1}_{0}(\Omega)$, and $H^{2}=H^{2}(\Omega)\cap H^{1}_{0}(\Omega)$. The operators $(-\Delta)^{s}$ are self-adjoints, i.e.,
$\left((-\Delta)^{s}u,v\right)=\left(u,(-\Delta)^{s}v\right),\quad\forall u,v\in H^{2s}.$
With the above notations we can define Hilbert spaces $\mathcal{H}_{s}:=H^{s+1}\times H^{s}$, endowed with the usual norms
[…]
In particular, with respect to the phase space corresponding to the weak solutions of problem (LABEL:P), the natural choice is
$\mathcal{H}:=\mathcal{H}_{0}=H^{1}\times H^{0}\equiv H^{1}_{0}(\Omega)\times L^{2}(\Omega),\\$
with norm given by
$U||^{2}_{\mathcal{H}}:=\|\nabla u\|^{2}+\|v\|^{2},\quad U\equiv(u,v).$
We assume the following hypotheses regarding the functions $h$, $g$ and $f$.
Assumption 2.1.
$h\in H^{0}$ is an external force and $f,g$ are given functions satisfying the following conditions:
  • $g\in C^{1}(\mathbb{R})$ is an increasing function such that $g(0)=0$, and there exist two positive constants $\kappa_{0}$ and $\kappa_{1}$ such that
    $\kappa_{0}|s|^{4}\leq g^{\prime}(s)\leq\kappa_{1}\left(1+|s|^{4}\right),\quad\forall s\in\mathbb{R}.$ (2.1)
  • $f\in C^{2}(\mathbb{R})$ with $f(0)=0$ is of the following growth condition: there exists a positive constant $L_{f}>0$ such that
    $\left|f^{\prime\prime}(s)\right|\leq L_{f}\left(1+|s|^{3}\right),\quad\forall s\in\mathbb{R},$ (2.2)
    in conjunction with the dissipative condition
    $\lim_{|s|\rightarrow\infty}\frac{f(s)}{s}>-\lambda_{1}$ (2.3)
    where $\lambda_{1}$ is the first eigenvalue of the Laplacian operator $\Delta$ with Dirichlet boundary condition.
Remark 2.1.
From Mean Value Theorem and (2.2), we get
$|f^{\prime}(s)|\leq C_{f}(1+|s|^{4}),\quad\forall s\in\mathbb{R},$ (2.4)
for some constant $C_{f}>0$. The dissipative condition (2.3) implies that there exists a constant $\nu\in[0,\lambda_{1})$ and $C_{\nu}>0$ such that
$-C_{\nu}-\frac{\nu}{2}|s|^{2}\leq F(s)\leq f(s)s+\frac{\nu}{2}|s|^{2},\quad\forall s\in\mathbb{R},$ (2.5)
where ${F(s):=\int_{0}^{s}f(\tau)d\tau}$.
Throughout the article we will use $0<\omega\leq 1$ as the constant defined by
$\omega:=1-\frac{\nu}{\lambda_{1}}>0.$ (2.6)
Both sources and the dissipation are referred to as critical in 3 dimensions. The criticality of the Sobolev’s embedding $H^{1}\hookrightarrow L^{6}(\Omega)$ implies the loss of compactness of the potential energy ${\int_{\Omega}F(u(x))dx}$ along with the loss of compactness of the damping $g(u_{t})$ from $H^{1}\rightarrow H^{-1}$.
2.2 Statements of the Main Results
Of main interest to this paper are the so called “weak solutions”. There are several notions of “weak solutions” which are often refereed to as “finite energy solutions”. However, in the critical case one needs to be careful regarding precise meaning of weak solutions. While proving existence of strong [regular solutions] is rather routine, existence and the properties of weak solutions are more elusive in the critical case. Typical notion would be the so called “generalized” solutions which are defined as strong limits of semigroup regular solutions. Under certain conditions such solutions can be shown to satisfy variational form of equations-the so called “weak solutions”. Weak solutions are not necessarily limits of strong solutions [unless some compactness related properties are assumed]- so these are of interest to the study in “critical” situations. Here, by critical we refer to the property that potential energy is just $L^{1}$ integrable without any compactness properties assumed in prior literature [16, 27] and references therein. In what follows, we shall adopt rather classical definition -following [5, Definition 2.1] where solutions to (LABEL:P) under the Assumption 2.1 were treated.
Definition 2.1.
For any $T>0$, a function $U(t)=(u(t),u_{t}(t))\in C_{w}(0,T;\mathcal{H})$, such that $u_{t}\in L^{6}(Q)$ and possessing the property $U(0)=(u(0),u_{t}(0))=(u_{0},u_{1})\in\mathcal{H}$ is said to be a weak energy solution to problem (LABEL:P) on interval $[0,T]$ iff Eq. (LABEL:P) is satisfied in the following variational sense:
$\begin{aligned} \int_{0}^{T}\int_{\Omega}\left[\,-u_{t}\psi_{t}+\nabla u\nabla\psi+g(u_{t})\psi+f(u)\psi\,\right]dxdt=-\left.\int_{\Omega}u_{t}\psi dx\right|_{0}^{T}+\int_{0}^{T}\int_{\Omega}h\psi dxdt,\end{aligned}$ (2.7)
holds any $\psi\in C(0,T;H^{1})\cap C^{1}(0,T;H^{0})\cap L^{6}(Q)$.
Here $Q:=\Omega\times(0,T)$ and $C_{w}(0,T,Y)$ denotes the space of weakly continuous functions with values in a Banach space $Y$.
The first result refers to an existence and properties of weak solutions as defined in Definition 2.1.
Theorem 2.2.
In reference to problem (LABEL:P), under Assumption 2.1, the following statements hold.
  • (I)
    Weak solutions: If initial data $U_{0}\in\mathcal{H}$, then problem (LABEL:P) has a weak solution $U=(u,u_{t})$ in the sense of Definition 2.1.
  • (II)
    Energy equality: The weak solutions satisfy the energy equality
    $E(U(t))+\int_{s}^{t}\int_{\Omega}g(u_{t})u_{t}dxds=E(U(s)),\quad\mbox{for all}\quad t\geq s\geq 0,$ (2.8)
    where $E(U)$ is the energy functional associated with problem (LABEL:P) given by the formula
[…]
  • (III)
    Hadamard well-posedness and stability: Let $U^{1}=(u^{1},u^{1}_{t})$ and $U^{2}=(u^{2},u^{2}_{t})$ be two weak solutions of problem (LABEL:P) with $U^{i}(0)=(u^{i}_{0},u^{i}_{1})=U^{i}_{0}\in\mathcal{H}$, $i=1,2$. Then
[…]
Remark 2.3.
Since weak solutions are constructed via Galerkin approximations, the uniqueness statement in Theorem 2.2 applied to weak solutions implies that all weak solutions can be constructed as Galerkin’s approximations.
The proof of the well-posedness theorem is based on the method introduced in [5]-which takes advantage of the enhanced dissipation exhibited by the additional integrability of the velocity. Existence of solutions was obtained in [7] by constructing suitable approximations of critical nonlinearities followed by the use of monotone operator theory. Uniqueness was shown in [5] by establishing first energy identity. However, continuous dependence stated in [5] holds for smoother initial data only. Since our ultimate goal is to have a properly defined dynamical system, we shall revisit the proof by using different method-based on Galerkin approximations. We will be able to establish full Hadamard well-posedness valid for all weak solutions. This is the main step toward construction of dynamical system $(\mathcal{H},S_{t})$, whose existence is guaranteed by Theorem 2.2. In addition, we shall show that all weak solutions can be obtained as strong limits of Galerkin solutions. This is in line with [20] and allows to use Galerkin approximation in the estimates where there is not enough of smoothness to use operational calculus and classical limiting arguments.
Models with nonlinear subquintic terms and enhanced dissipation were considered earlier in [16, 27]. However, in the case of weak solutions uniqueness of solutions and Hadamard well-posedness was considered an open problem. This issue has been positively resolved in [5] and in the present paper.
The main results of this work describe long-time behavior of solutions generated by the dynamical system $(\mathcal{H},S_{t})$. The definitions of the respectful quantities along with the known properties are relegated, for the convenience of a reader, to the Appendix 7.
Theorem 2.4.
[Global attractor] Assume that the hypotheses of Theorem 2.2 hold. Then, the associated dynamical system $(\mathcal{H},S_{t})$ corresponding to the dynamics in (LABEL:P) has a compact global attractor $\mathfrak{A}$ of the form $\mathfrak{A}=\mathrm{M}^{u}(\mathcal{N})$, where $\mathrm{M}^{u}(\mathcal{N})$ is the unstable manifold emanating from the set of stationary solution $\mathcal{N}$. Moreover, $\mathfrak{A}$ consists of full trajectories $\gamma=\{S_{t}U_{0}=U(t):t\in\mathbb{R}\}$ such that
$\lim_{t\rightarrow-\infty}\mbox{dist}_{\mathcal{H}}(U(t),\mathcal{N})=0\quad\mbox{and}\quad\lim_{t\rightarrow+\infty}\mbox{dist}_{\mathcal{H}}(U(t),\mathcal{N})=0.$
The proof of the existence of a global attractor is based on an adaptation of “Ball’s” method that employs energy arguments and is fundamentally based on the validity of energy equality valid for weak solutions. This method is adapted to the present context by incorporating a nonlinear dissipation. The latter is possible due to monotonicity properties.
Under the additional assumptions imposed on the damping $g^{\prime}(0)>0$, the following properties of the said attractor $\mathfrak{A}$ are valid.
Theorem 2.5.
[Quasi-stability property] Let the Assumption (2.1) holds and $g^{\prime}(0)>0$. Then the dynamical system $(\mathcal{H},S_{t})$ is quasi-stable.
The proof that the dynamical system is quasi-stable follows from the fact that the dynamical system $(\mathcal{H},S_{t})$ satisfies the so called quasi-stability inequality, which is an enhanced version of squeezing property -see the Appendix 7.
Theorem 2.6.
[Properties of the attractor] Under the Assumptions of Theorem 2.5 the following properties of the attractor $\mathfrak{A}$ are valid.
  • (I)
    Finite-dimensionality: the global attractor $\mathfrak{A}$ has finite fractal dimension $dim^{\mathcal{H}}_{f}\mathfrak{A}$;
  • (II)
    Regularity: any full trajectory $\gamma=\{U(t):t\in\mathbb{R}\}$ from attractor $\mathfrak{A}$ enjoys the following regularity properties
    $U=(u,u_{t})\in L^{\infty}(\mathbb{R};H^{2}\times H^{1}),\quad u_{tt}\in L^{\infty}(\mathbb{R};H^{0}).$
    Moreover, there exists $R_{*}>0$ such that
    $\|u_{tt}(t)\|^{2}+\|u_{t}(t)\|^{2}_{H^{1}}+\|u(t)\|^{2}_{H^{2}}\leq R^{2}_{*},$
    for all $t\in\mathbb{R}$ and some positive constant $R_{*}$.
  • (III)
    Exponential attractor: the dynamical system $(\mathcal{H},S_{t})$ possesses a generalized fractal exponential attractor $\mathfrak{A}_{\exp}$ with finite dimension in the extended space $\mathcal{H}_{-s}:=H^{-s+1}\times H^{-s}$ with $0<s\leq 1$.
The proof of Theorem 2.6 is based on the validity of the quasi-stability property of the system $(\mathcal{H},S_{t})$. This property is related to an exponential rate of approximation of sets that attract each other, therefore, the assumption that $g^{\prime}(0)>0$ is expected.
The result presented above are inspired by earlier papers [5, 7] where the model with critical [quintic] sources and enhanced dissipations were considered. For these models an existence and uniqueness of weak solutions has been established in references cited above. Hadamard well-posedness has been shown in [5] for subcritical sources only. Questions related to blow up of weak solutions have been investigated in [4, 6, 8, 15]. Full Hadamard well-posedness and complete analysis of long time dynamics is the main contribution of the present manuscript.
We conclude this section by briefly outlining general approach taken in the paper. The existence of weak solutions is established by Galerkin method with suitable $L^{p}$ global estimates. Weak solutions are shown to satisfy energy equality. Here, again, enhanced dissipation plays a major role. This is a critical step in order to prove uniqueness of weak solutions and Hadamard well-posedness. The latter allows to construct a well-posed dynamical system. By resorting, again, to enhanced dissipation, the system $(\mathcal{H},S_{t})$ is shown to be ultimately dissipative, hence admitting weak global attractor. The next critical step is asymptotic compactness. Relying on already proved energy identity, an adaptation of the method of [1] to account for nonlinear damping, allows to establish asymptotic compactness, hence an existence of global attractor. Under the additional assumption of non-degeneracy of the damping at the origin, global attractor is shown to be: smooth and also finite-dimensional. In addition, the dynamical system has also exponential attractor. These last properties follow from the established quasi-stability inequality.
2.3 Past Contributions.
Wave dynamics in 3 dimensions with subquintic and quintic source has been an object of long-time and intense activities. The main issues of interest to this work are well-posedness and long time behavior. The presence of the quintic source makes the traditional methodology un-applicable-due to the lack of appropriate embeddings which would “close” the needed estimates. Relatively recent development of Strichartz estimates opened the door to novel methodologies, where the additional space integrability is achieved due to compensation with time integrability. Careful balance of time-space integrability allows to construct local solutions referred to as Shatah-Struwe [S-S] solutions. However, such solutions do not admit global extensions [unless the case is subcritical]. Here, the rescue comes by appealing to non-concentration argument [9]. This argument relies on finite speed of propagation, hence sensitive to nonlinear damping. And this is one of the main points where nonlinearity of the damping in the present model becomes predicament. In fact, in the case of linearly bounded dissipation, the methodology pursued in [20]. In the case of critical sources and linear damping is the most effective. The starting point of the analysis in [20] is local existence of S-S-solutions. These are obtained by using Strichartz estimates applied to Galerkin solutions. Since Strichartz estimates do not admit global extensions [unless the case is subcritical], global existence of Galerkin limits is obtained by appealing to non-concentration argument [9]. Having established global existence [S-estimates do not blow up in time]-the next step is a construction of a dynamical system which contains weak Galerkin limits [not necessarily S-solutions which may be lost in the limiting process]. The latter requires uniqueness and continuous dependence of the said solutions. In order to achieve this, the method of backward trajectories is employed in order to generate smoothness of backward trajectories. Classical forward regularity argument allows to propagate the smoothness to any time $T$. This allows to claim uniqueness of orbits and also energy equality on any time interval. Equipped with the above properties, energy method is used in [20] in order to assert an asymptotic compactness and to conclude the proof of existence of global attractor. The orbits in the attractor become eventually S-S solutions.
In the case of subquintic sources, the analysis is simpler [S-S solutions are global], though still technical. Some of the recent relevant references include [24, 28] and references within, where nonlinear dissipation has been treated. For the quintic waves with nonlinear dissipation some of the elements of the S-S method described above has been very recently applied to subquintic and quintic waves with nonlinear enhanced dissipation [30]. In particular, backward trajectories method patterned after [29] is the driving argument in establishing asymptotic compactness of the trajectories and their smoothness. This requires strong assumptions imposed on the damping [nondegenearcy of the slope at the origin] and also the source [coercivity of quintic estimate].
One oof the goals of this paper is to show that such restrictions, in general, are not needed. The model with enhanced dissipation takes its own life with no reference to either S-S solutions or backward trajectories. The main protagonist is the enhanced dissipation which could also be degenerate $g^{\prime}(0)=0$.] This alone provides asymptotic compactness effect for a large range of sources including the quintic. No coercivity assumption is needed. The role of Strichartz estimates is taken, in some sense, by the enhanced dissipation generated by the nonlinearity of the damping. However, in the present case one is in a position to show that the enhanced dissipation is retained in the ultimate dissipativity estimate [unlike Strichartz estimate]. In view of this, the present work is a continuation of the model considered [5] in the direction of dynamical systems and its long time behaviour. The main player in this game is energy equality established for weak solutions - an earlier version of the equality can be found in [5]. Adaptation of J.Ball’s method to nonlinear dissipation leads to the existence of global attractor. A general quasi-stability property for the dynamics is established under additional assumption of nonvanishing derivative of the damper. The latter implies finite-dimensionality, enhanced regularity of the attractor and also an existence of exponential attractor. In order to recover maximal regularity of the attractor, the splitting method used by Chueshov and Lasiecka in [12] in conjunction with gradient structure of the attractor and properties of the stationary set. In conclusion, the results provided eliminate critical restrictions in [30] and moreover lead to different methodology applicable to many other hyperbolic like dynamics with critical sources. The method is much simpler and it could also be applied to nonlocal damping under certain assumptions [30, 31].
One should also mention works where strong damping is added to the model. But this is a form of regularization where originally hyperbolic dynamics become parabolic [with linearisation generating an analytic semigroup]. The latter is a very different problem. References [10, 11, 17, 19].
The reminder of this paper is devoted to the proofs of the three theorems. Appendix 7 at the end of the paper, collects various definitions and well known facts used in the manuscript.
3 Proof of Theorem 2.2: Weak Solutions
Here is the plan of attack. Weak solutions will be constructed via taking the limits of Galerkin approximations. The next step is establishing energy identity valid for each weak solution. This is done via special discretization method introduced by H. Koch [21]. The latter allows to establish uniqueness of weak solutions. This step also allows to conclude that each weak solution can be obtained via Galerkin limit. Hadamard well-posedness is obtained via the estimates performed on Galerkin approximations and their limits.
3.1 Proof of Theorem 2.2-(I): Construction of Weak Solutions
The proof of existence of weak solutions proceeds along general strategy used in [5], where the enhanced dissipation plays critical role in the arguments. However, the construction in [5] is based on suitable truncations of nonlinear terms and heavy use of monotone operator theory [3, 22]. In the present case, we resort to Galerkin solutions, which also give a nice computational framework. The advantage of Galerkin solutions is that “formal” computations are well justified and lead to conserved quantities. However, critical limit passages on nonlinear terms rely on some arguments seen already in [5]. On the other hand, [5] deals also with the boundary sources of critical range.These will be more difficult to incorporate within the Galerkin scheme.
Galerkin’s approximate solutions
Let $0<\lambda_{1}\leq\lambda_{2}\leq\cdots$ be the eigenvalues of the Laplacian operator $\Delta$ with Dirichlet boundary condition and $\omega_{1},\omega_{2},\cdots$ be the corresponding eigenfunctions such that they form an orthonormal basis in $H^{0}$. We assume the boundary $\Gamma=\partial\Omega$ regular enough such that $\omega_{j}\in H^{2}.$ Let $V_{k}=\mbox{Span}\{\omega_{1},\cdots,\omega_{k}\}$ the subspace of $H^{2}$ generated by the first $k$ elements of $\{\omega_{j}\}_{j\in\mathbb{N}}$. For $k\in\mathbb{N}$ we can construct a function $u^{k}$ given by
$u^{k}(t)=\sum_{j=1}^{k}y_{jk}(t)\,\omega_{j}\in V_{k},\quad t\in[0,t_{k}),$
where $(y_{jk})$ is a local solution on $0,t_{k})\in[0,T)$ of the following system of ODEs:
$\left\{\begin{array}[]{l}\displaystyle{(u^{k}_{tt}(t),\omega_{j})+(\nabla u^{k}(t),\nabla\omega_{j})+(g(u^{k}_{t}(t)),\omega_{j})+(f(u^{k}(t)),\omega_{j})=(h,\omega_{j})}\\ \displaystyle{U^{k}(0)=(u^{k}(0),u^{k}_{t}(0))=(u_{0k},u_{1k})=U_{0k},\quad j=1,\cdots,k.}\end{array}\right$. (3.3)
In what follows, it is necessary to establish a priori estimates that allow the extension of the local solution to the entire interval $[0,T]$. With this estimate in hand, we can then take the limit in the approximate problem (3.3) to establish the existence of weak solutions to (LABEL:P).
A priori estimates
We consider the approximate system (3.3) with
$U^{k}(0)\longrightarrow U_{0}\quad\mbox{strongly in}\quad\mathcal{H}.$ (3.4)
Taking $\omega_{j}=u^{k}_{t}(t)$ in the approximate equation (3.3) and integrating over $[0,t]$ yields
$E(U^{k}(t))+\int_{0}^{t}g(u^{k}_{t})u^{k}_{t}dxds=E(U^{k}(0)),$ (3.5)
where $E$ is given in (2.9). Next, from (2.5) and embedding $H^{1}\hookrightarrow H^{0}$, we get
$\int_{\Omega}F(u^{k})dx\geq-\frac{\nu}{2}\|u^{k}(t)\|^{2}-C_{\nu}|\Omega|\geq-\frac{\nu}{2\lambda_{1}}\|\nabla u^{k}(t)\|^{2}-C_{\nu}|\Omega|.$ (3.6)
And, from Hölder inequality, embedding $H^{1}\hookrightarrow H^{0}$, and Young’s inequality, we have
$\int_{\Omega}hu^{k}dx\leq\|h\|\|u^{k}(t)\|\leq\frac{1}{\lambda_{1}^{1/2}}\|h\|\|\nabla u^{k}(t)\|\leq\frac{1}{\delta\lambda_{1}}\|h\|^{2}+\frac{\delta}{4}\|\nabla u^{k}(t)\|^{2},$ (3.7)
for any $\delta>0$. Then, choosing in particular $\delta=\omega$ (as defined in (2.6)), combining estimates (3.6) and (3.7), we can estimate the energy $E(U^{k})$ lower as follows
[…]
Using (2.1), (3.8), and convergence (3.4), it follows from (3.5) that
[…]
in $0,t_{k})$, $\forall k\in\mathbb{N}$. Now let’s estimate the term $E(U_{0k})$. Indeed, from (2.4)-(2.5), we have
$\int_{\Omega}F(u_{0k})dx$ $\leq$ $\int_{\Omega}f(u_{0k})u_{0k}dx+\frac{\nu}{2}\|u_{0k}\|^{2}$
$=$ $\int_{\Omega}\int_{0}^{1}f^{\prime}(\theta u_{0k})d\theta u_{0k}^{2}dx+\frac{\nu}{2}\|u_{0k}\|^{2}$
$\leq$ $C_{f}\int_{\Omega}\int_{0}^{1}(1+|\theta u_{0k}|^{4})d\theta u_{0k}^{2}dx+\frac{\nu}{2}\|u_{0k}\|^{2}$
$\leq$ $\left(C_{f}+\frac{\nu}{2}\right)\|u_{0k}\|^{2}+C_{f}\|u_{0k}\|^{6}_{6}.$
Hence, using embeddings $H^{1}\hookrightarrow L^{6}(\Omega)\hookrightarrow H^{0}$ and convergence (3.4), we have
$E(U_{0\kappa})\leq\frac{1}{2}\|u_{1k}\|^{2}+\frac{1}{2}\|\nabla u_{0k}\|^{2}+\left(C_{f}+\frac{\nu}{2}\right)\|u_{0k}\|^{2}+C_{f}\|u_{0k}\|^{6}_{6}+\|h\|\|u_{0k}\|\leq C,$
[…]
for $t\in[0,t_{k})$. Therefore, estimate (3.10) allows us to extend the local solution of the approximate problem to whole interval $[0,T]$, for any given $T>0$. From (3.10), we have
$U^{k}=(u^{k},u^{k}_{t})\quad$ is bounded in $\quad L^{\infty}(0,T;\mathcal{H}),$ (3.11)
$(u^{k}_{t})\quad$ is bounded in $\quad L^{6}(Q).$ (3.12)
Limiting process.
Now, we proceed to take the limit in the approximate problem (3.3), thereby demonstrating that a limiting solution satisfies the variational problem (2.7).
Analysis of linear terms.
From (3.11), we extract a subsequence still denoted by itself such that
$U^{k}\rightharpoonup U=(u,v)\quad\mbox{weakly* in}\quad L^{\infty}(0,T;\mathcal{H}).$ (3.13)
It is easy to see that $v=u_{t}$ and convergence (3.13) allows passing the limit in the linear terms of the approximated Eq. (3.3).
Analysis of the source term $f(u^{k})$.
From (3.11), Applying Lions’ compactness theorem [23, Ch. 1, Th. 5.1 ] with $H^{1}\hookrightarrow\hookrightarrow H^{0}\hookrightarrow H^{0}$, it turns out that there exists a subsequence of $(u^{k})$ still denoted by $(u^{k})$ such that
$u^{k}\rightarrow u\quad\mbox{strongly in}\quad H^{0}.$
Passing to a subsequence if necessary, we get
$u^{k}\rightarrow u\quad\mbox{a.e. in}\quad Q.$ (3.14)
Using the continuity of $f$, it follows from (3.14) that
$f(u^{k})\rightarrow f(u)\quad\mbox{a.e. in}\quad Q.$ (3.15)
Now, from Assumption (2.4), embeddings $H^{1}\hookrightarrow L^{6}(\Omega)\hookrightarrow L^{6/5}(\Omega)$, and (3.11), we have
$\int_{0}^{T}\int_{\Omega}|f(u^{k})|^{6/5}dxdt\leq C\int_{0}^{T}\int_{\Omega}[\,|u^{k}|^{6/5}+|u^{k}|^{6}\,]dxdt<\infty.$
That is,
$f(u^{k})\quad\mbox{in bounded in}\quad L^{6/5}(Q).$ (3.16)
Then, using (3.15) and (3.16), it follows from Lions’ lemma [23, Ch. 1, Lemma 1.3] that
$f(u^{k})\rightarrow f(u)\quad\mbox{weakly in}\quad L^{6/5}(Q)=[L^{6}(Q)]^{\prime}.$ (3.17)
Thus, from (3.17), we have
$\int_{0}^{T}(f(u^{k}),\psi)dt\longrightarrow\int_{0}^{T}(f(u),\psi)dt,\quad\mbox{as}\quad k\rightarrow+\infty,\quad\forall\psi\in L^{6}(Q).$
Analysis of the damping term $g(u^{k}_{t})$.
From Assumption (2.1), embeddings $L^{6}(\Omega)\hookrightarrow H^{0}\hookrightarrow L^{6/5}(\Omega)$, and (3.12), we have
$\int_{0}^{T}\int_{\Omega}|g(u^{k}_{t})|^{6/5}dxds\leq C\int_{0}^{T}\int_{\Omega}[\,|u^{k}_{t}|^{6/5}+|u^{k}_{t}|^{6}]dxds<\infty.$
That is,
$g(u^{k}_{t})\quad\mbox{is bounded in}\quad L^{6/5}(Q).$
Consequently, we get
$g(u^{k}_{t})\rightarrow\chi\quad\mbox{weakly in}\quad L^{6/5}(Q),\quad\mbox{for some}\quad\chi\in L^{6/5}(Q).$
From Eq. (3.3), we also have $u^{k}_{tt}\rightarrow u_{tt}$ weakly* in $L^{\infty}(0,T;H^{-1})$. Therefore, letting $k\rightarrow+\infty$ in Eq. (3.3) and making use of above convergence, we conclude that
$u_{tt}-\Delta u+\chi+f(u)=h\quad\mbox{in}\quad H^{-1}\quad\mbox{for a.e.}\;\;t\in[0,T].$
It remains to show that $g(u_{t})=\chi$. To prove the latter we shall use maximal monotonicity of the operator $g(u_{t})$ along with [3, Lemma 2.3]. From this result, it suffices to show that:
$\lim\sup_{k,n\rightarrow+\infty}\int_{0}^{T}\int_{\Omega}\left[\,g(u_{t}^{k})-g(u_{t}^{n})\,\right]\left[\,u_{t}^{k}-u_{t}^{n}\,\right]dxdt\leq 0,$ (3.18)
where the above inequality is expected to hold on Galerkin solutions $u^{n},u^{k}$. Considering the difference $z^{k,m}=u^{k}-u^{n}$ in (3.3) and taking the multiplier $z^{k,n}_{t}$, the following equality emerges
[…]
Integrating (3.19) from 0 to $T$, we obtain
[…]
Due to the convergence of the initial data (3.4) the equality (3.20) above implies
$\lim\sup_{k,n\rightarrow+\infty}\int_{0}^{T}\int_{\Omega}\left[g(u_{t}^{k})-g(u_{t}^{n})\right]z^{k,n}_{t}dxdt\leq\lim\sup_{k,n\rightarrow+\infty}\int_{0}^{T}\int_{\Omega}\left[f(u^{k})-f(u^{n})\right]z_{k,n_{t}}dxdt.$
If $p<5$, then the last term in the above inequality is zero. However, for $p=5$ more delicate argument is needed. which depends on enhanced dissipation. To this end, we rewrite
$\begin{aligned} \int_{\Omega}\left[f(u^{k})-f(u^{n})\,\right]z^{k,n}_{t}dx=&\;\int_{\Omega}\int_{0}^{1}f^{\prime}(\theta u^{k}+(1-\theta)u^{n})d\theta z^{k,n}z^{k,n}_{t}dx\\ =&\;\int_{\Omega}\int_{0}^{1}f^{\prime}(\theta u^{k}+(1-\theta)u^{n})d\theta\frac{1}{2}\frac{d}{dt}|z^{k,n}|^{2}dx\\ =&\;\frac{1}{2}\frac{d}{dt}\left[\int_{\Omega}\int_{0}^{1}f^{\prime}(\theta u^{k}+(1-\theta)u^{n})d\theta|z^{k,n}|^{2}dx\right]\\ &-\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\theta u^{k}+(1-\theta)u^{n})\left(\,\theta u^{k}_{t}+(1-\theta)u^{n}_{t}\,\right)d\theta|z^{k,n}|^{2}dx.\end{aligned}$ (3.21)
Then, substituting (3.21) in (3.19), we have
[…]
$+\int_{\Omega}\left[g(u_{t}^{k})-g(u_{t}^{n})\right]z^{k,n}_{t}dx=\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\theta u^{k}+(1-\theta)u^{n})\left(\,\theta u^{k}_{t}+(1-\theta)u^{n}_{t}\,\right)d\theta|z^{k,n}|^{2}dx.$
From the dissipativity condition (2.3), there exist $\mu\in[0,\lambda_{1})$ and $N>0$ large such that
$f^{\prime}(s)\geq-\mu,\quad\mbox{for}\quad|s|>N.$ (3.23)
On the other hand, from (2.4), we have
$|f^{\prime}(s)|\leq C_{f}(1+|N|^{4}),\quad\mbox{for}\quad|s|\leq N.$ (3.24)
So, setting $\Omega=\Omega_{1}\cup\Omega_{2}$, where
$\Omega_{1}:=\{x\in\Omega:|\theta u^{k}+(1-\theta)u^{n}|>N\}\quad\mbox{and}\quad\Omega_{2}:=\{x\in\Omega:|\theta u^{k}+(1-\theta)u^{n}|\leq N\},$
it follows from (3.23) and (3.24) that
$\begin{aligned} &\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime}(\theta u^{k}+(1-\theta)u^{n})d\theta|z^{k,n}|^{2}dx\\ &=\frac{1}{2}\left[\int_{\Omega_{1}}+\int_{\Omega_{2}}\right]\int_{0}^{1}f^{\prime}(\theta u^{k}+(1-\theta)u^{n})d\theta|z^{k,n}|^{2}dx\\ &\geq-\frac{\mu}{2}\int_{\Omega_{1}}|z^{k,n}|^{2}dx-\frac{C_{f}(1+|N|^{4})}{2}\int_{\Omega_{2}}|z^{k,n}|^{2}dx\geq-\frac{K_{f}}{2}\int_{\Omega}|z^{k,n}|^{2}dx,\end{aligned}$ (3.25)
where $K_{f}=\max\{\mu,C_{f}(1+|N|^{4})\}$.
On the other hand, from (2.4), Hölder inequality with $\frac{2}{3}+\frac{1}{3}=1$, embedding $H^{1}\hookrightarrow L^{6}(\Omega)$, and (3.11), we have
[…]
Thus, setting
[…]
it follows from (3.25) and (3.26) that
[…]
Now, using the definition of the functional $E_{z^{k,n}}$, we can rewrite (3.22) as follows
$\begin{aligned} &\frac{d}{dt}E_{z^{k,n}}(T)+\int_{\Omega}\left[g(u_{t}^{k})-g(u_{t}^{n})\right]|z^{k,n}|^{2}dx\\ &=\underbrace{K_{f}\int_{\Omega}z^{k,n}z^{k,n}_{t}dx}_{I_{1}}+\underbrace{\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\theta u^{k}+(1-\theta)u^{n})\left(\,\theta u^{k}_{t}+(1-\theta)u^{n}_{t}\right)d\theta|z^{k,n}|^{2}dx}_{I_{2}}.\end{aligned}$ (3.28)
From embedding $H^{1}\hookrightarrow H^{0}$, and (3.27), it is easy to see that
$\left|I_{1}\right|\leq\frac{K_{f}}{\lambda_{1}^{1/2}}\|z^{k,n}_{t}(t)\|\|\nabla z^{k,n}(t)\|\leq\frac{K_{f}}{2\lambda_{1}^{1/2}}\left[\|z^{k,n}_{t}(t)\|^{2}+\|\nabla z^{k,n}(t)\|^{2}\right]\leq\frac{K_{f}}{\lambda_{1}^{1/2}}E_{z^{k,n}}(t).$ (3.29)
From (2.2), Hölder inequality with $\frac{1}{2}+\frac{1}{6}+\frac{1}{3}=1$, embedding $H^{1}\hookrightarrow L^{6}(\Omega)$, (3.27), and (3.11), we have
$\begin{aligned} |I_{2}|\leq&\;C\left[\,1+\|u^{k}(t)\|_{6}^{3}+\|u^{m}(t)\|_{6}^{3}\,\right]\left[\,\|u^{k}_{t}(t)\|_{6}+\|u^{n}_{t}(t)\|_{6}\,\right]\|\|z^{n,k}(t)\|_{6}^{2}\\ \leq&\;C\psi(t,u^{k},u^{n})E_{z^{k,n}}(t),\end{aligned}$ (3.30)
where $\psi(t,u^{k},u^{n}):=\|u^{k}_{t}(t)\|_{6}^{6}+\|u^{n}_{t}(t)\|_{6}^{6}\in L^{1}(0,T)$.
Thus, using (3.29) and (3.30), returning to (3.28), we obtain
$\frac{d}{dt}E_{z^{k,n}}(t)+\int_{\Omega}\left[g(u_{t}^{k})-g(u_{t}^{n})\right]z^{k,n}_{t}dx\leq C\psi(t,u^{k},u^{n})E_{z^{k,n}}(t).$ (3.31)
Using the multiplier $e^{-C\int_{0}^{t}\psi(s,u^{k},u^{n})ds}$ and integrating (3.31) from 0 to $t$, we conclude that
$\begin{aligned} \underbrace{E_{z^{k,n}}(t)}_{\geq 0}+\int_{0}^{t}\int_{\Omega}\underbrace{e^{C\int_{s}^{t}\psi(\tau,u^{k},u^{n})d\tau}}_{\geq 1}\left[g(u_{t}^{k})-g(u_{t}^{n})\right]z^{k,n}_{t}dxds\leq e^{C\int_{0}^{t}\psi(s,u^{k},u^{n})ds}E_{z^{k,n}}(0),\end{aligned}$ (3.32)
for all $t\in[0,T]$. Hence, using (3.12), (3.27), and (3.4), it follows from (3.32) that
[…]
for all $t\in[0,T]$. Therefore, (3.33) implies the validity of (3.18), which by virtue of [3, Lemma 2.3] provides the identification of the weak limit.
Therefore, letting $k\rightarrow+\infty$ in (3.3) and making use of above convergence show that $U=(u,u_{t})$ is a weak solution claimed in Theorem 2.2 satisfying
$u_{tt}-\Delta u+g(u_{t})+f(u)=h\quad$
in the sense of Definition 2.1. This concludes the proof of Part $\bf(I)$} of Theorem 2.2.
Remark 3.1.
We note that the same conclusion as above holds under weaker assumptions imposed on the source $f$. It suffices that the constraints on the lower bound of $f;$ is imposed only for critical level of nonlinearity: $\lim_{|s|\rightarrow\infty}\frac{f^{\prime}(s)}{s^{4}}>-\nu$. The subcritical terms will converge to zero by the virtue of the compactness.
3.2 Proof of Theorem 2.2-(II): Energy Equality
Energy equality [as opposed to inequality] is a critical tool in asserting well-posedness of the dynamical system generated by weak solutions of the system. The latter are typically shown to be weakly continuous, a consequence of apriori bounds. Energy equality allows an upgrade to continuity. In addition, energy identity is also fundamental in proving asymptotic compactness. The latter is particularly important when dealing with nonlinearities of critical nature, where standard methods based on squeezing estimates are noneffective. Instead, long time horizon estimates of a given solution along the trajectory provides more flexibility. In order to take advantage of this technique [1], energy equality is essential.
While energy identity can be formally obtained by a simple integration by parts with a suitable multiplier, the real issue is a justification of this procedure in the absence of sufficient regularity of weak solutions. In the case of the model under study, one would like to multiply the equation by just velocity $u_{t}$. However, $L^{p}$ membership of this element does not permit formal calculations. The definition of weak solutions requires testing with $H^{1}$ elements. It may be noted that energy inequality cen be derived via weak lower semicontinuity from the weak limits of approximated solutions. However, the method will not produce equality. In order to comply with this requirement, we shall follow procedure originated in [21] where time derivatives of solutions are suitably approximated by rather special finite difference method. Thus, to derive the energy identity, we employ a finite difference approximation of the time derivatives, following the approach used in [2, Appendix 5].
To proceed- let $v\in B(0,T;V)$ where $V$ is a Hilbert space and $B([0,T];V)$ is the space of $V$-value functions which are bounded on $[0,T]$, endowed with usual norm $\|y\|_{B([0,T];V)}=\sup_{t\in[0,T]}\|y(t)\|_{V}$. Let $\epsilon>0$ be a small parameter destined to go to zero and consider the extension $v(t)$ to all $t\in\mathbb{R}$ defined by
$v(t)=\left\{\begin{array}[]{l}v(T)\quad\mbox{if}\quad t\geq T,\\ v(0)\quad\mbox{if}\quad t\leq 0.\end{array}\right$.
With the above extensions we define
$v^{+}_{\epsilon}(t)$ $=$ $v(t+\epsilon)-v(t),$
$v^{-}_{\epsilon}(t)$ $=$ $v(t)-v(t-\epsilon),$
$D_{\epsilon}v(t)$ $=$ $\frac{1}{2\epsilon}\left[v^{+}_{\epsilon}(t)+v^{-}_{\epsilon}(t)\right].$
The following proposition stated in [5, Proposition 1] is a slight extension of the corresponding result proved in [2, Proposition 5.2] and [21] where it was used within the context of nonlinear dynamic elasticity.
Proposition 3.2.
  • Let $v$ be a weakly continuous with the values in a Hilbert space $V$. Then
    $\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\left(v(t),D_{\epsilon}v(t)\right)_{V}dt=\frac{1}{2}\left[\,\|v(T)\|^{2}_{V}-\|v(0)\|^{2}_{V}\,\right].$ (3.35)
  • For $v\in W^{1,p}(0,T;W)$, the following limits are well defined in $L^{p}(0,T;W)$, for all $1<p<\infty$ and any Banach space $W$:
    $\lim_{\epsilon\rightarrow 0}D_{\epsilon}v=v_{t},\quad\lim_{\epsilon\rightarrow 0}\frac{1}{\epsilon}v^{+}_{\epsilon}=v_{t},\quad\lim_{\epsilon\rightarrow 0}\frac{1}{\epsilon}v^{-}_{\epsilon}=v_{t}.$ (3.36)
    Moreover, if $v_{t}$ is weakly continuous with the values in $W$, then for every $t\in(0,T)$, $D_{\epsilon}v(t)\rightarrow v_{t}(t)$ weakly in $W$ and
    $\lim_{\epsilon\rightarrow 0}v^{-}_{\epsilon}(T)=v_{t}(T),\quad\lim_{\epsilon\rightarrow 0}\frac{1}{\epsilon}v^{+}_{\epsilon}(0)=v_{t}(0),\quad\mbox{weakly in}\quad W.$
  • Let $H\subset V\subset H^{\prime}$, $v_{tt}\in L^{2}(0,T;H^{\prime})$ and $v\in L^{2}(0,T;H)$. Then
    $\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\left(v_{tt}(t),D_{\epsilon}v(t)\right)_{V}dt=\frac{1}{2}\left[\,\|v_{t}(T)\|^{2}_{V}-\|v_{t}(0)\|^{2}_{V}\,\right].$ (3.37)
Next, using variational form (2.7) of problem (LABEL:P) with $\psi=D_{\epsilon}u(t)\in H^{1}$ leads to:
$\begin{aligned} &\int_{0}^{T}u_{tt}D_{\epsilon}udxdt+\int_{0}^{T}\int_{\Omega}\nabla u\nabla D_{\epsilon}udxdt\\ &+\int_{0}^{T}\int_{\Omega}g(u_{t})D_{\epsilon}udxdt+\int_{0}^{T}\int_{\Omega}f(u)D_{\epsilon}udxddt=\int_{0}^{T}\int_{\Omega}hD_{\epsilon}udxdt.\end{aligned}$ (3.38)
By direct application of (3.37), we obtain:
$\lim_{\epsilon\rightarrow 0}\int_{0}^{T}u_{tt}D_{\epsilon}udxdt=\frac{1}{2}\|u_{t}(T)\|^{2}-\frac{1}{2}\|u_{t}(0)\|^{2}.$ (3.39)
And, by direct application of (3.35), we get
$\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\int_{\Omega}\nabla uD_{\epsilon}(\nabla u)dxdt=\frac{1}{2}\|\nabla u(T)\|^{2}-\frac{1}{2}\|\nabla u(0)\|^{2}.$ (3.40)
Using Assumption (2.1) and that $u_{t}\in L^{6}(Q)$, we have
$\int_{0}^{T}\int_{\Omega}|g(u_{t})|^{6/5}dxdt\leq C\int_{0}^{T}\int_{\Omega}[\,|u_{t}|^{6/5}+|u_{t}|^{6}\,]dxdt<\infty.$
That is, $g(u_{t})\in L^{6/5}(Q).$ On the other hand, using $u_{t}\in L^{6}(Q)$, it follows from (3.36) that
$\lim_{\epsilon\rightarrow 0}D_{\epsilon}(u)=u_{t}\in L^{6}(Q).$
Hence,
$\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\int_{\Omega}g(u_{t})D_{\epsilon}udxdt=\int_{0}^{T}\int_{\Omega}g(u_{t})u_{t}dxdt.$ (3.41)
Now, using Assumption (2.4), embedding $H^{1}\hookrightarrow L^{6}(\Omega)$ and that $u(t)\in H^{1}$ for all $t\in[0,T]$, we get
$\int_{0}^{T}\int_{\Omega}|f(u)|^{6/5}dxdt\leq C\int_{0}^{T}\int_{\Omega}[\,|u|^{6/5}+|u|^{6}\,]dxdt<\infty.$
That is, $f(u)\in L^{6/5}(Q).$ Again using $\lim_{\epsilon\rightarrow 0}D_{\epsilon}(u(t))=u_{t}\in L^{6}(Q),$ we conclude that
$\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\int_{\Omega}f(u_{t})D_{\epsilon}udxddt=\int_{0}^{T}\int_{\Omega}f(u)u_{t}dxddt.$ (3.42)
Finally, it follows directly from (3.36) that
$\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\int_{\Omega}hD_{\epsilon}udxddt=\int_{0}^{T}\int_{\Omega}hu_{t}dxddt.$ (3.43)
Thus, taking the limit as $\epsilon\rightarrow 0$ in (3.38) and using convergences (3.39)-(3.43), we conclude that
$\begin{aligned} &\frac{1}{2}\left[\,\|u_{t}(T)\|^{2}-\|u_{t}(0)\|^{2}\,\right]+\frac{1}{2}\left[\,\|\nabla u(T)\|^{2}-\|\nabla u(0)\|^{2}\,\right]\\ &+\int_{0}^{T}\int_{\Omega}g(u_{t})u_{t}dxddt+\int_{0}^{T}\int_{\Omega}f(u)u_{t}dxdt=\int_{0}^{T}\int_{\Omega}hu_{t}dxdt.\end{aligned}$ (3.44)
Therefore, using that
$\int_{0}^{T}\int_{\Omega}f(u)u_{t}dxdt=\int_{\Omega}F(u(T))dx-\int_{\Omega}F(u(0))dx$
and
$\int_{0}^{T}\int_{\Omega}hu_{t}dxdt=\int_{\Omega}hu(T)dx-\int_{\Omega}hu(0)dx,$
we obtain from (3.44) the desired energy identity for weak solutions (2.8). This concludes the proof of Part $\bf(II)$}.
3.3 Proof of Theorem 2.2-(III): Hadamard Well-posedness
Hadamard wellposedness of weak solutions includes two additional properties of the solutions: uniqueness and continuous dependence on the data [often refereed to as stability or robustness]. In fact, the key property is the uniqueness. Having energy identity and the uniqueness of weak solutions allows to deduce the robustness [21]. The latter needs to be shown for the already constructed solutions - and not necessarily from the definition of weak solutions [due to uniqueness]. Let $u^{1}(t)\neq u^{2}(t)$ be two different weak solutions of problem (LABEL:P) with $(u^{1}(0),u^{1}_{t}(0))=(u^{1}_{0},u^{1}_{1})\in\mathcal{H}$ and $(u^{2}(0),u^{2}_{t}(0))=(u^{2}_{0},u^{2}_{1})\in\mathcal{H}$. Using the energy identity (2.8) for the difference $z(t):=u^{1}(t)-u^{2}(t)$, we obtain the following equality valid for all weak solutions
[…]
From Assumption (2.1), we obtain
$\begin{aligned} \int_{\Omega}\left[g(u^{1}_{t})-g(u^{2}_{t})\right]z_{t}dx=&\;\int_{\Omega}\int_{0}^{1}g^{\prime}(\theta u^{1}_{t}+(1-\theta)u^{2}_{t})d\theta|z_{t}|^{2}dx\\ \geq&\;\kappa_{0}\int_{\Omega}\int_{0}^{1}|\theta u^{1}_{t}+(1-\theta)u^{2}_{t}|^{4}d\theta|z_{t}|^{2}dx\geq 0.\end{aligned}$ (3.46)
On the other hand, as in (3.21) we can rewrite
$\begin{aligned} \int_{\Omega}\left[f(u^{1})-f(u^{2})\,\right]z_{t}dx=&\;\frac{1}{2}\frac{d}{dt}\left[\int_{\Omega}\int_{0}^{1}f^{\prime}(\theta u^{1}+(1-\theta)u^{2})d\theta|z|^{2}dx\right]\\ &-\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\theta u^{1}+(1-\theta)u^{2})\left(\,\theta u^{1}_{t}+(1-\theta)u^{2}_{t}\,\right)d\theta|z|^{2}dx.\end{aligned}$ (3.47)
Then, using (3.46) and (3.47), it follows from (3.45) that
$\begin{aligned} \frac{d}{dt}E_{z}(t)\leq\underbrace{K_{f}\int_{\Omega}zz_{t}dx}_{\mathrm{I}_{1}}+\underbrace{\frac{1}{2}\int_{\Omega}\int_{0}^{1}\left|f^{\prime\prime}(\theta u^{1}+(1-\theta)u^{2})\right|\left|\,\theta u^{1}_{t}+(1-\theta)u^{2}_{t}\,\right|d\theta|z|^{2}dx}_{\mathrm{I}_{2}},\end{aligned}$ (3.48)
where
[…]
with the constant $K_{f}$ obtained as in (3.25). By estimates analogous to (3.25) and (3.26) we obtain the following result
[…]
And, as in (3.29) and (3.30) we also obtain
$\left|\mathrm{I}_{1}\right|\leq\frac{K_{f}}{\lambda_{1}^{1/2}}E_{z}(t)\quad\mbox{and}\quad\left|\mathrm{I}_{2}\right|\leq C\psi(t,u^{1},u^{2})E_{z}(t),$
where $\psi(t,u^{1},u^{2})=1+\|u^{1}_{t}(t)\|_{6}^{6}+\|u^{2}_{t}(t)\|_{6}^{6}$. Substituting $|\mathrm{I}_{1}|$ and $|\mathrm{I}_{2}|$ into (3.48), we have
$\frac{d}{dt}E_{z}(t)\leq C\psi(t,u^{1},u^{2})E_{z}(t).$ (3.50)
Applying Gronwall’s lemma in (3.50), we obtain
$E_{z}(t)\leq e^{C\int_{0}^{t}\psi(s,u^{1},u^{2})ds}E_{z}(0),\quad\forall t\in[0,T].$ (3.51)
Therefore, using equivalence (3.49) and (3.12), it follows from (3.51) that
[…]
Remark 3.3.
  1. 1.
    The arguments given above provide uniqueness and continuous dependence in “one shot”. This should be contrasted with [5] where continuous dependence requires additional regularity of solutions.
  2. 2.
    The proof of the Theorem above shows that every weak solution can be obtained as a strong limit of Galerkin solutions. This remark is useful when one attempts to provide smooth approximations of weak solutions.
  3. 3.
    Energy identity allows to upgrade in standard manner weak continuity of solutions to strong continuity removing $C_{w}$ from the characterization of the solutions. As a consequence, solutions form well defined dynamical system.
3.4 Generation of a Dynamical System
Recall that a dynamical system is a pair of objects $(X,S_{t})$ consisting of a complet metric space $X$ and a family of continuous mappings $\{S_{t}:\,t\in\mathbb{R}^{+}\}$ of $X$ into itself with the semigroup properties:
$S_{0}=I,\quad S_{t+\tau}=S_{t}\circ S_{\tau}.$
$S_{t}$ is called evolution semigroup (or evolution operator). Moreover, $\{S_{t}:\,t\in\mathbb{R}^{+}\}$ is a strogly continuous semigroup of bounded linear operators on $X$ if
$\lim_{t\rightarrow 0}S_{t}x=x\quad\mbox{for every}\quad x\in X.$
It is also called: semigroup of class $C_{0}$ or simply a $C_{0}$-semigroup.
Theorem 2.2 guarantees that weak solutions to problem (LABEL:P) generate a dynamical system $(\mathcal{H},S_{t})$. The evolution operator $S_{t}:\mathcal{H}\to\mathcal{H}$ is characterized by the relation
$S_{t}U_{0}=(u(t),u_{t}(t))=U(t),\quad U_{0}\in\mathcal{H},$
where $U$ denotes the unique weak solution to equation (LABEL:P). Moreover, it directly follows from Theorem 2.2-$\bf(III)$} that the dynamical system $(\mathcal{H},S_{t})$ exhibits the local Lipschitz continuity property on every interval $[0,T]$.
$\|S_{t}U_{0}^{1}-S_{t}U_{0}^{2}\|_{\mathcal{H}}^{2}\leq e^{c_{R}t}\|U_{0}^{1}-U_{0}^{2}\|_{\mathcal{H}}^{2},\quad\forall t\in[0,T],$ (3.52)
where $U_{0}^{i}\in\mathcal{H}$ with $\|U^{i}_{0}\|_{\mathcal{H}}\leq R$, $i=1,2$ and $c_{R}$ is a positive constant depending on $R$.
4 Proof of Theorem 2.4: Global Attractors
The proof of the existence of a compact global tractor is established by the results obtained in Propositions 4.1 and 4.3 below. Proposition 4.1 shows that the dynamical system $(\mathcal{H},S_{t})$ is ultimately dissipative and Proposition 4.3 shows that $(\mathcal{H},S_{t})$ is asymptotically compact. Therefore, the existence of a compact global attractor follows as a direct application of Theorem 7.1 (in Appendix 7). Furthermore, the characterization of the attractor $\mathfrak{A}\equiv\mathcal{M}^{u}(\mathcal{N})$ is guaranteed by applying Theorem 7.2 using Proposition 4.4, demonstrated below, which guarantees that the dynamic system $(\mathcal{H},S_{t})$ is gradient. The details follow.
4.1 Ultimate Dissipativity
Proposition 4.1.
Under the assumptions of Theorem 2.2, the dynamical system $(\mathcal{H},S_{t})$ has an absorbing set $\mathcal{B}$ in $\mathcal{H}$.
Proof.
Let $B$ be a bounded subset in $\mathcal{H}$ and let $U(t)=S_{t}U_{0}$ be a weak solution of problem (LABEL:P) with $U_{0}\in\mathcal{H}$. In order, let us consider the modified functional
$\mathcal{J}(U(t)):=E(U(t))+L,$
where $L:=\frac{1}{\omega\lambda_{1}}\|h\|^{2}+C_{\nu}|\Omega|$ and $E$ is the energy functional defined in (2.9). By argument similar to those used in (3.8), we obtain
[…]
Using the energy equality (2.8) with $s=t$ and $t=t+1$, we obtain
$\int_{t}^{t+1}\int_{\Omega}g(u_{t})u_{t}dxds=E(U(t))-E(U(t+1))=\mathcal{J}(U(t))-\mathcal{J}(U(t+1))=:\mathcal{D}(t).$
From Assumption (2.1), we get
$\int_{t}^{t+1}\|u_{t}(s)\|^{6}_{6}ds\leq\frac{1}{\kappa_{0}}\mathcal{D}(t).$ (4.2)
From Hölder inequality with $\frac{2}{3}+\frac{1}{3}=1$, it follows from (4.2) that
$\int_{t}^{t+1}\|u_{t}(s)\|^{2}ds\leq|\Omega|^{2/3}\left(\int_{t}^{t+1}\|u_{t}(s)\|^{6}_{6}ds\right)^{1/3}\leq\frac{|\Omega|^{2/3}}{\kappa_{0}^{1/3}}\mathcal{D}(t)^{1/3}.$ (4.3)
Furthermore, using Mean Value Theorem (MVT) and (4.3) there exist $t_{1}\in[t,t+\frac{1}{4}]$ and $t_{2}\in[t+\frac{3}{4},t+1]$ such that
$\|u_{t}(t_{j})\|^{2}\leq 4\int_{t}^{t+1}\|u_{t}(s)\|^{2}ds\leq\frac{4|\Omega|^{2/3}}{\kappa_{0}^{1/3}}\mathcal{D}(t)^{1/3},\quad j=1,2.$ (4.4)
Next, multiplying Eq. (LABEL:P) by $u(t)$ and integrating over $t_{1},t_{2}]\times\Omega$, we have
$\begin{aligned} \int_{t_{1}}^{t_{2}}E(s)ds+\frac{1}{2}\int_{t_{1}}^{t_{2}}\|\nabla u(s)\|^{2}ds=&\;\frac{3}{2}\int_{t_{1}}^{t_{2}}\|u_{t}(s)\|^{2}ds-\int_{t_{1}}^{t_{2}}\int_{\Omega}g(u_{t})udxds\\ &\;+\int_{\Omega}u(t_{1})u_{t}(t_{1})dx-\int_{\Omega}u(t_{2})u_{t}(t_{2})dx\\ &\;+\int_{t_{1}}^{t_{2}}\left[\int_{\Omega}F(u)dx-\int_{\Omega}f(u)udx\right]ds.\end{aligned}$ (4.5)
From Assumption (2.5), we get
$\int_{t_{1}}^{t_{2}}\left[\int_{\Omega}F(u)dx-\int_{\Omega}f(u)udx\,\right]ds\leq\frac{\nu}{2}\int_{t_{1}}^{t_{2}}\int_{\Omega}|u|^{2}dxds\leq\frac{\nu}{2\lambda_{1}}\int_{t_{1}}^{t_{2}}\|\nabla u(s)\|^{2}ds.$ (4.6)
Then, taking the sum of $L$ on both sides of (4.5) and, using the definition of $\mathcal{J}(U(t))$ and (4.6), we get
$\begin{aligned} \int_{t_{1}}^{t_{2}}\mathcal{J}(U(s))ds+\frac{\omega}{2}\int_{t_{1}}^{t_{2}}\|\nabla u(s)\|^{2}ds\leq&\;\frac{3}{2}\int_{t_{1}}^{t_{2}}\|u_{t}(s)\|^{2}ds-\int_{t_{1}}^{t_{2}}\int_{\Omega}g(u_{t})udxds\\ &\;+\int_{\Omega}u(t_{1})u_{t}(t_{1})dx-\int_{\Omega}u(t_{2})u_{t}(t_{2})dx+L.\end{aligned}$ (4.7)
Using Assumption (2.1), Hölder and Young inequalities, embeddings $H^{1}\hookrightarrow L^{6}(\Omega)\hookrightarrow H^{0}$, (4.2), and (4.3), we have
$-\int_{t_{1}}^{t_{2}}\int_{\Omega}g(u_{t})udxds$ $\leq$ $\kappa_{1}\int_{t_{1}}^{t_{2}}\int_{\Omega}(1+|u_{t}|^{4})|u_{t}||u|dxds$
$\leq$ $\kappa_{1}\left(\int_{t}^{t+1}\|u_{t}(s)\|^{2}ds\right)^{1/2}\left(\int_{t}^{t+1}\|u(s)\|^{2}ds\right)^{1/2}$
$+\kappa_{1}\left(\int_{t_{1}}^{t_{2}}\|u_{t}(s)\|^{6}_{6}ds\right)^{5/6}\left(\int_{t}^{t+1}\|u(s)\|^{6}_{6}ds\right)^{1/6}$
$\leq$ $\frac{\kappa_{1}}{\lambda_{1}^{1/2}}\left(\int_{t}^{t+1}\|u_{t}(s)\|^{2}ds\right)^{1/2}\left(\int_{t}^{t+1}\|\nabla u(s)\|^{2}ds\right)^{1/2}$
$+\frac{\kappa_{1}}{\lambda^{1/2}}\left(\int_{t_{1}}^{t_{2}}\|u_{t}(s)\|^{6}_{6}ds\right)^{5/6}\left(\int_{t}^{t+1}\|\nabla u(s)\|^{6}ds\right)^{1/6}$
$\leq$ $\frac{2\kappa_{1}}{(\omega\lambda_{1})^{1/2}}\left(\frac{|\Omega|^{2/3}}{\kappa_{0}^{1/3}}\mathcal{D}(t)^{1/3}\right)^{1/2}\sup_{t\leq s\leq t+1}\mathcal{J}(U(s))^{1/2}$
$+\frac{2\kappa_{1}}{(\omega\lambda)^{1/2}}\left(\frac{1}{\kappa_{0}}\mathcal{D}(t)\right)^{5/6}\sup_{t\leq s\leq t+1}\mathcal{J}(U(s))^{1/2}$
$\leq$ $\frac{2\kappa_{1}^{2}|\Omega|^{2/3}}{\varepsilon\omega\lambda_{1}\kappa_{0}^{1/3}}\mathcal{D}(t)^{1/3}+\frac{2\kappa_{1}}{\varepsilon\omega\lambda\kappa_{0}^{5/3}}\mathcal{D}(t)^{5/3}+\varepsilon\sup_{t\leq s\leq t+1}\mathcal{J}(U(s)).$
From embedding $H^{1}\hookrightarrow H^{0}$ and (4.4), we get
$\int_{\Omega}u(t_{1})u_{t}(t_{1})dx-\int_{\Omega}u(t_{2})u_{t}(t_{2})dx$ $\leq$ $\|u(t_{1})\|\|u_{t}(t_{1})\|+\|u(t_{2})\|\|u_{t}(t_{2})\|$
$\leq$ $\frac{1}{\lambda_{1}^{1/2}}\left[\,\|\nabla u(t_{1})\|\|u_{t}(t_{1})\|+\|\nabla u(t_{2})\|\|u_{t}(t_{2})\|\,\right]$
$\leq$ $\frac{4|\Omega|^{1/3}}{\kappa_{0}^{1/6}\lambda_{1}^{1/2}}\mathcal{D}(t)^{1/6}\sup_{t\leq s\leq t+1}\|\nabla u(s)\|$
$\leq$ $\frac{16|\Omega|^{2/3}}{\varepsilon\omega\kappa_{0}^{1/3}\lambda_{1}}\mathcal{D}(t)^{1/3}+\varepsilon\sup_{t\leq s\leq t+1}\mathcal{J}(U(t)).$
Returning to (4.5), we obtain
$\int_{t_{1}}^{t_{2}}\mathcal{J}(U(s))ds$ $\leq$ $C_{0}\mathcal{D}(t)^{1/3}+C_{1}\mathcal{D}(t)^{5/3}+2\varepsilon\sup_{t\leq s\leq t+1}\mathcal{J}(U(s))+L,$ (4.8)
where $C_{0}:=\frac{|\omega|^{2/3}}{\kappa_{0}^{1/3}}+\frac{2\kappa_{1}^{2}|\Omega|^{2/3}}{\varepsilon\omega\lambda_{1}\kappa_{0}^{1/3}}+\frac{16|\Omega|^{2/3}}{\varepsilon\omega\kappa_{0}^{1/3}\lambda_{1}}$, $C_{1}=\frac{2\kappa_{1}}{\varepsilon\omega\lambda\kappa_{0}^{5/3}}$.
Using that $\mathcal{J}$ is decreasing and MVT for integrals, there exists $\tau\in[t_{1},t_{2}]$ such that
$\mathcal{J}(U(t+1))\leq\mathcal{J}(U(\tau))=\frac{1}{t_{2}-t_{1}}\int_{t_{1}}^{t_{2}}\mathcal{J}(U(s))ds\leq 2\int_{t_{1}}^{t_{2}}\mathcal{J}(U(s))ds.$ (4.9)
Hence, from (4.2), (4.9), and (4.8), we get
$\mathcal{J}(U(t))$ $=$ $\mathcal{J}(U(t+1))+\mathcal{D}(t)\leq 2\int_{t_{1}}^{t_{2}}\mathcal{J}(U(s))ds+\mathcal{D}(t)$
$\leq$ $2C_{0}\mathcal{D}(t)^{1/3}+2C_{1}\mathcal{D}(t)^{5/3}+\mathcal{D}(t)+4\varepsilon\sup_{t\leq s\leq t+1}\mathcal{J}(U(s))+2L.$
Choosing $\varepsilon=\frac{1}{8}$ and using $\mathcal{J}(U(t))=\sup_{t\leq s\leq t+1}\mathcal{J}(U(s))$, we obtain
$\mathcal{J}(U(t))\leq\left[\,4C_{0}+4C_{1}\mathcal{D}(t)^{4/3}+2\mathcal{D}(t)^{2/3}\,\right]\mathcal{D}(t)^{1/3}+4L.$
From the last inequality and using that $(a+b)^{3}\leq 4(a^{3}+b^{3})$, we have
$\mathcal{J}(U(t))^{3}\leq 4\left[\,4C_{0}+4C_{1}\mathcal{D}(t)^{4/3}+2\mathcal{D}(t)^{2/3}\,\right]^{3}\left[\,\mathcal{J}(U(t))-\mathcal{J}(U(t+1))\,\right]+[2^{8/3}L]^{3}.$ (4.10)
Using the definition of $\mathcal{D}(t)$ and the energy equality (2.8) we obtain that $|\mathcal{D}(t)|\leq 2|E(U_{0})|$. Then, we infer that
$C_{0}:=\sup_{U_{0}\in B}4\left[\,4C_{0}+4C_{1}\mathcal{D}(t)^{4/3}+2\mathcal{D}(t)^{2/3}\,\right]^{3}<\infty.$
Thus, (4.10) can be rewritten as
$\mathcal{J}(U(t))^{3}\leq C_{0}\left[\,\mathcal{J}(U(t))-\mathcal{J}(U(t+1))\,\right]+\left[2^{8/3}L\right]^{3}.$ (4.11)
Applying Nakao’s lemma [26, Lemma 2.1] to (4.11), we conclude that
$\mathcal{J}(U(t))\leq\left(2C_{0}^{-1}(t-1)^{+}+\mathcal{J}(U(0))^{-2}\right)^{-\frac{1}{2}}+2^{8/3}L.$ (4.12)
Using (4.1), it follows from (4.12) that
[…]
and
[…]
Then, choosing $R^{2}>\frac{2^{14/3}L}{\omega}$, we conclude, from (4.13) and (4.14) that the ball
[…]
is an absorbing set in $\mathcal{H}$, for the dynamical system $(\mathcal{H},S_{t})$. ∎
Remark 4.2.
In the particular case, when $h\equiv 0$ and $C_{\nu}=0$ in (2.5) we have that $\mathcal{J}(U(t))=E(U(t))$ and $L=0$. Thus, in (4.14) we obtain that
[…]
That is, the dynamical system $(\mathcal{H},S_{t})$ possesses, in its natural topological phase space $\mathcal{H}$ a compact global attractor $\mathfrak{A}=\{(0,0)\in\mathcal{H}\}$ (the set of a single point).
4.2 Asymptotic Compactness
Proposition 4.3.
Let the assumptions of Theorem 2.2 hold. Then, the dynamical system $(\mathcal{H},S_{t})$ associated with problem (LABEL:P) is asymptotically compact.
Proof.
From Theorem 2.2-$\bf(IV)$} every weak solution $U(t)=(u(t),u_{t}(t))$ of Eq. (LABEL:P) satisfies the energy equality:
$E(U(t))+\int_{0}^{t}\int_{\Omega}g(u_{t})u_{t}dxds=E(U(0)),$ (4.15)
where
[…]
Consider the standard perturbed Lyapunov’s function $V$ with a perturbation $\epsilon$ given by
$V(U):=E(U)+\epsilon\int_{\Omega}u_{t}udx.$
Since identity (2.8) is valid for weak solutions of problem (LABEL:P) and perturbation is given by compact term, standard argument [using the equation] leads to the energy identity satisfied by each weak solution in the the perturbed functional.
$V(U(T))=e^{-\epsilon T}V(U(0))+\int_{0}^{T}e^{-\epsilon(T-t)}H(U(t))dt,$ (4.16)
where $H(U)=L(U)+M(U)$ with
[…]
$M(U)$ $:=$ $-\epsilon\left[\int_{\Omega}f(u)udx-\int_{\Omega}F(u)dx\right]-\frac{\epsilon}{2}\|\nabla u\|^{2}-\epsilon\int_{\Omega}g(u_{t})udx+\epsilon^{2}\int_{\omega}u_{t}udx.$
Our goal is to show that a semiflow $\mathcal{G}$- family of maps $\varphi:[0,\infty)\rightarrow\mathcal{H}$ [weak solutions] is asymptotically compact. We shall follow closely [including the notation] the method presented in [1] applied to linear dissipation. We take a sequence $\varphi_{j}\in\mathcal{G}$ with $\varphi_{j}(0)$ bounded, and let $t_{j}\rightarrow\infty$. By (2.8) $E(\varphi_{j}(t_{j}))$ is bounded which also implies $\varphi_{j}(t_{j})$ is bounded in $\mathcal{H}$ . This allows to select weakly convergent subsequence
$\varphi_{j}(t_{j})\rightharpoonup\lambda,\quad\lambda\in\mathcal{H}$ (4.17)
and by continuity of the semiflow, there exists $\lambda_{T}\in\mathcal{H}$ such that
$\varphi_{j}(t_{j}-T)\rightharpoonup\lambda_{T},\quad\lambda_{T}\in\mathcal{H}.$
Using [1, (C4w)] we can still assume that there exists a trajectory $\widetilde{\varphi}(t)=(u(t),u_{t}(t))\in\mathcal{G}$ with
$U^{j}(t)\equiv\varphi_{j}(t_{j}+t-T)\rightharpoonup\widetilde{\varphi}(t)=(u(t),u_{t}(t))=U(t)\in\mathcal{G},$
where $\widetilde{\varphi}(0)=\lambda_{T}$ and $\widetilde{\varphi}(T)=\lambda.$ Applying the perturbed energy equality (4.16) to the above solutions by substituting $U^{j}(t)\equiv\varphi_{j}(t_{j}+t-T)$ we obtain
$V(\varphi_{j}(t_{j}))=e^{-\epsilon T}V(\varphi_{j}(t_{j}-T)+\int_{0}^{T}e^{-\epsilon(T-t)}H(U^{j}(t))dt$
and the equality for limiting solutions [using again energy equality in (4.16)]
$V(\lambda)=V(\widetilde{\varphi}(T))=e^{-\epsilon T}V(\lambda_{T})+\int_{0}^{T}e^{-\epsilon(T-t)}H(U(t))dt.$
Our goal is to show that
$\lim\sup_{j\rightarrow\infty}V(\varphi_{j}(t_{j}))\leq V(\lambda)\leq\lim\inf_{j\rightarrow\infty}V(\varphi_{j}(t_{j})).$ (4.18)
The second inequality follows from weak lower semicontinuity of $V(\varphi)$ and weak convergence in (4.17). Also, in the case when $H$ is compact, the inequality in (4.18) follows quickly by taking $T$ to infinity. In non-compact case, more work is needed as noted in [1]. The key observation is the control of “sign” which allows to use version of Fatou’s lemma. We shall use the elementary relation: $\lim\sup(F)=-\lim\inf(-F).$ The main task is to pass with the weak limit under the integral sign. For transparency, we shall focus on critical-noncompact terms, where passage with the weak limit requires additional [to weak convergence] properties. In short, the last two terms in $M(U)$ owning to compact embedding $H^{1}(\Omega)\subset L^{5}(\Omega)$ satisfy
$\lim_{j}\left[\epsilon\int_{\Omega}g(u_{t}^{j})u^{j}dx+\epsilon^{2}\int_{\Omega}u_{t}^{j}u^{j}dx\right]=\epsilon\int_{\Omega}g(u_{t})udx+\epsilon^{2}\int_{\Omega}u_{t}udx.$
Since ${-\epsilon\left[\int_{\Omega}f(u^{j})u^{j}dx-\int_{\Omega}F(u^{j})dx\right]-\frac{\epsilon}{2}\|\nabla u^{j}\|^{2}}$ is negative, weak lower semicontinuity implies
$\int_{0}^{T}e^{-\epsilon(T-t)}\left[\epsilon\int_{\Omega}(f(u)u-F(u))dx+\frac{\epsilon}{2}\|\nabla u(t)\|^{2}\right]dt$
$\leq\lim\inf_{j}\int_{0}^{T}e^{-\epsilon(T-t)}\left[\epsilon\int_{\Omega}(f(u^{j})u^{j}-F(u^{j}))dx+\frac{\epsilon}{2}\|\nabla u^{j}(t)\|^{2}\right]dt$
$=-\lim\sup_{j}\int_{0}^{T}e^{-\epsilon(T-t)}\left[-\epsilon\int_{\Omega}(f(u^{j})u^{j}-F(u^{j}))dx-\frac{\epsilon}{2}\|\nabla u^{j}(t)\|^{2}\right]dt.$
This gives
$-\int_{0}^{T}e^{-\epsilon(T-t)}M(U)dt$ $\leq$ $\lim\inf_{j}\int_{0}^{T}e^{-\epsilon(T-t)}(-M(U^{j}))dt$
$=$ $-\lim\sup_{j}\int_{0}^{T}e^{-\epsilon(T-t)}M(U^{j})dt.$
Hence
$\lim\sup_{j}\int_{0}^{T}e^{-\epsilon(T-t)}M(U^{j})dt\leq\int_{0}^{T}e^{-\epsilon(T-t)}M(U)dt.$ (4.19)
We shall next estimate the damping terms: For any small constant $\eta>0$
[…]
$\int_{0}^{T}e^{-\epsilon(T-t)}L(U)dt+\epsilon C_{\eta}\leq\lim\inf_{j}\int_{0}^{T}e^{-\epsilon(T-t)}[-L(U^{j})+\epsilon C_{\eta}]dt$
and converting into lim sup
$\int_{0}^{T}e^{-\epsilon(T-t)}L(U)dt\geq\lim\sup_{j}\int_{0}^{T}e^{-\epsilon(T-t)}L(U^{j})dt.$ (4.20)
Combining (4.19) and (4.20), we get
$\lim\sup_{j}\int_{0}^{T}e^{-\epsilon(T-t)}\left[\,L(U^{j})+M(U^{j})\,\right]dt\leq\int_{0}^{T}e^{-\epsilon(T-t)}\left[\,L(U)+M(U)\,\right]dt.$
This yields
$\lim\sup_{j}\int_{0}^{T}e^{-\epsilon(T-t)}H(U^{j})dt\leq\int_{0}^{T}e^{-\epsilon(T-t)}H(U)dt.$
By taking $T\rightarrow\infty$ and using boundedness of trajectories in $\mathcal{G}$ one obtains the conclusion in (4.18). The final result follows from the fact that weak convergence and norm convergence imply strong convergence. ∎
4.3 Gradient System
Proposition 4.4.
Assume that the assumptions of Theorem 2.2 hold. Then, $\varphi:\mathcal{H}\to\mathbb{R}$ given by
[…]
is a strict Lyapunov functional for the dynamical system $(\mathcal{H},S_{t})$. Consequently, $(\mathcal{H},S_{t})$ is a gradient dynamical system.
Proof.
Let us define $\varphi:=E$. From energy equality (2.8), we have
$\varphi(S_{t}U_{0})+\int_{0}^{t}\int_{\Omega}g(u_{t})u_{t}dxd\tau=\varphi(U_{0}),$ (4.21)
for every $U_{0}\in\mathcal{H}$. From Assumption (2.1), we have
$\int_{0}^{t}\int_{\Omega}g(u_{t})u_{t}dxd\tau\geq\kappa_{0}\int_{0}^{t}\int_{\Omega}|u_{t}|^{6}dxd\tau.$ (4.22)
Then, it follows from (4.21) and (4.22) that $\varphi(S_{t}U_{0})=\varphi(U_{0})$. Hence
$\int_{0}^{t}\|u_{t}(\tau)\|^{6}_{6}\,d\tau\leq 0.$
Using embedding $L^{6}(\Omega)\hookrightarrow H^{0}$, we conclude that
$\|u_{t}(t)\|=0\ \ \hbox{for almost everywhere}\ \ t>0.$
Moreover, it follows from $u_{t}\in C([0,T];H^{0})$ for all $T>0$ that $\|u_{t}(t)\|=0$ for all $t>0$, which implies that $U_{0}\in\mathcal{N}$, where $\mathcal{N}$ is the set of stationary points of the dynamical system $(\mathcal{H},S_{t})$. Using that
$\quad U_{0}\in\mathcal{N}\quad\Leftrightarrow\quad S_{t}(U_{0})=U_{0},\quad t>0,$
then $\varphi$ is a strict Lyapunov functional for the dynamical system $(\mathcal{H},S_{t})$. ∎
5 Proof of Theorem 2.5: Quasi-stability property
The existence of a compact global attractor $\mathfrak{A}$ is guaranteed by Theorem 2.4 under weaker assumption $g^{\prime}(0)\geq 0$. However, a strict inequality $g^{\prime}(0)>0$ allows to show that the dynamical system $(\mathcal{H},S_{t})$ is quasi-stable according to Definition 7.8. And within the context of the theory of dynamical systems associated with the study of attractors, for quasi-stable systems, it is possible to obtain important properties such as the finite dimension and smoothness of the attractor, as well as the existence of exponential attractors. Properties that will be established for the dynamical system associated with problem (LABEL:P) in Theorem 2.6.
The proof that the system $(\mathcal{H},S_{t})$ is quasi-stable follows from the fact that $(\mathcal{H},S_{t})$ satisfies conditions (7.1) and (7.2) of the Definition 7.8 of a quasi-stable system. The condition (7.1) is guaranteed directly from the Lipschitz property (3.52). It remains to prove the quasi-stability inequality (7.2) which is ensured by the relation (5.1) established in the following proposition.
5.1 Quasi-stability Estimate
Proposition 5.1.
Let $(u^{1}(t),u^{1}_{t}(t))=S_{t}(u^{1}_{0},u^{1}_{1})$ and $(u^{2}(t),u^{2}_{t}(t))=S_{t}(u^{2}_{0},u^{2}_{t})$ be two weak solutions to (LABEL:P) with initial data $U^{1}_{0}=(u^{1}_{0},u^{1}_{1}),U^{2}_{0}=(u^{2}_{0},u^{2}_{1})$ lying in a bounded set $B\subset\mathcal{H}$. Let the hypotheses of Theorem 2.2 be valid. In addition we assume that $g^{\prime}(0)>0$. Let $z(t)=u^{1}(t)-u^{2}(t).$ Then we have the following relation
[…]
where $b(t)$ and $c(t)$ are nonnegative scalar functions satisfying $b\in L^{1}(\mathbb{R}^{+})$ with $\lim_{t\rightarrow+\infty}b(t)=0$ and $c(t)$ is bounded on $0,\infty]$.
Proof.
We begin with the following simple observation: Using that $g\in C^{1}(\mathbb{R})$ and $g^{\prime}(0)>0$, then there exists a constant $\eta>0$ such that $g^{\prime}(s)>0$ for all $|s|<2\eta$. And since $g^{\prime}>0$ in $\mathbb{R}$, there exists a constant $m>0$ such that $g^{\prime}(s)\geq m>0$ for all $|s|\leq\eta$. On the other hand, from Assumption (2.1), we have $g^{\prime}(s)\geq\kappa_{0}|s|^{4}\geq\kappa_{0}\eta^{4}>0$ for all $|s|\geq\eta$. Hence, choosing $\kappa_{2}=\inf\{m,\kappa_{0}\eta^{4}\}$, we infer
$g^{\prime}(s)\geq\kappa_{2},\quad\mbox{for all}\quad s\in\mathbb{R}.$ (5.2)
Therefore, from (5.2) and using the same arguments used in [18, Remark 4:1] with $\gamma=4$, we obtain
$\left[\,g(r)-g(s)\,\right](r-s)\geq\kappa_{2}|s-r|^{2}+\frac{\kappa_{0}}{10}\left(|r|^{4}+|s|^{4}\right)|r-s|^{2}.$ (5.3)
Let $z=u^{1}-u^{2}.$ Then $z$ satisfies the equation
$\left\{\begin{array}[]{l}\displaystyle{z_{tt}-\Delta z+\left[\,g(u^{1}_{t})-g(u^{2}_{t})\,\right]+\left[\,f(u^{1})-f(u^{2})\,\right]=0,\quad\mbox{in}\quad\Omega\times[0,\infty),}\\ \displaystyle{(z(0),z_{t}(0))=(u_{0}-v_{0},u_{1}-v_{1}),\quad z=0,\quad\mbox{on}\quad\Gamma.}\end{array}\right$. (5.6)
Using the energy identity (2.8) for the difference of two weak solutions, we have
$\begin{aligned} \frac{d}{dt}E_{z}(t)+\int_{\Omega}D(z_{t})dx=&\;K_{f}\int_{\Omega}zz_{t}dx\\ &+\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\theta u^{1}+(1-\theta)u^{2})\left[\theta u^{1}_{t}+(1-\theta)u^{2}_{t}\right]d\theta|z|^{2}dx.\end{aligned}$ (5.7)
where
$D(z_{t}):=\left[g(u^{1}_{t})-g(u^{2}_{t})\right]z_{t}$
and
[…]
with the constant $K_{f}$ obtained from dissipativity condition (2.3) as in (3.25).
Integrating (5.7) from $\tau\geq 0$ to $T$, we get
$\begin{aligned} &E_{z}(T)-E_{z}(\tau)+\int_{\tau}^{T}\int_{\Omega}D(z_{t}(s))dxds\\ &=\underbrace{K_{f}\int_{\tau}^{T}\int_{\Omega}zz_{t}ds}_{J_{1}}+\underbrace{\frac{1}{2}\int_{\tau}^{T}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\theta u^{1}+(1-\theta)u^{2})\left[\theta u^{1}_{t}+(1-\theta)u^{2}_{t}\right]d\theta|z|^{2}dxds}_{J_{2}}.\end{aligned}$ (5.8)
Using Hölder inequality, (5.3), and Young’s inequality, we have
$|J_{1}|$ $\leq$ $K_{f}\int_{\tau}^{T}\|z\|\|z_{t}\|ds\leq\frac{K_{f}}{\kappa_{2}^{1/2}}\left(\int_{\tau}^{T}\|z\|^{2}ds\right)^{1/2}\left(\int_{\tau}^{T}\int_{\Omega}D(z_{t}(s))dxds\right)^{1/2}$
$\leq$ $\frac{K_{f}^{2}}{2\kappa_{2}}\int_{\tau}^{T}\|z\|^{2}ds+\frac{1}{2}\int_{\tau}^{T}\int_{\Omega}D(z_{t}(s))dxds.$
From Assumption (2.2), Hölder inequality with $\frac{1}{2}+\frac{1}{6}+\frac{1}{3}=1$, embedding $H^{1}\hookrightarrow L^{6}(\Omega)$, and Young’s inequality, we have
$|J_{2}|$ $\leq$ $\;C\int_{\tau}^{T}\left[\,1+\|u^{1}\|^{3}_{6}+\|u^{2}\|^{3}_{6}\,\right]\left[\,\|u^{1}_{t}\|_{6}+\|u^{2}_{t}\|_{6}\,\right]\|z\|_{6}^{2}ds$
$\leq$ $C_{B}\int_{\tau}^{T}\left[\,\|u^{1}_{t}\|_{6}+\|u^{2}_{t}\|_{6}\,\right]\|\nabla z\|^{2}ds$
$\leq$ $\delta\int_{\tau}^{T}E_{z}(s)ds+C_{B,\delta}\int_{\tau}^{T}d(s,u^{1},u^{2})E_{z}(s)ds,$
where
$d(t,u^{1},u^{2}):=\|u^{1}_{t}(t)\|^{6}_{6}+\|u^{2}_{t}(t)\|^{6}_{6}.$
Substituting $|J_{1}|$ and $|J_{2}|$ into (LABEL:ASP-11), we find
$\begin{aligned} E_{z}(T)-E_{z}(\tau)+\frac{1}{2}\int_{\tau}^{T}\int_{\Omega}D(z_{t}(s))dxds\leq&\;\frac{K_{f}^{2}}{2\kappa_{2}}\int_{\tau}^{T}\|z\|^{2}ds+\delta\int_{\tau}^{T}E_{z}(s)ds\\ &+C_{B,\delta}\int_{\tau}^{T}d(s,u^{1},u^{2})E_{z}(s)ds.\end{aligned}$ (5.9)
Now, integrating (5.9) from 0 to $T$, we obtain
$\begin{aligned} TE_{z}(T)+\frac{1}{2}\int_{0}^{T}\int_{\tau}^{T}\int_{\Omega}D(z_{t}(s))dxdsd\tau\leq&\;\int_{0}^{T}E(\tau)d\tau+\frac{K_{f}^{2}}{2\kappa_{2}}\int_{0}^{T}\int_{\tau}^{T}\|z\|^{2}dsd\tau\\ &+\delta T\int_{0}^{T}E_{z}(s)ds+TC_{B,\delta}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds.\end{aligned}$ (5.10)
Next, multiplying Eq. (5.6) by $z$ and integrating over $\Omega\times[0,T]$, we have
$\begin{aligned} 2\int_{0}^{T}E_{z}(s)ds=&\;2\int_{0}^{T}\|z_{t}\|^{2}ds+K_{f}\int_{0}^{T}\|z\|^{2}ds\\ &\;-\underbrace{\left[\int_{\Omega}z_{t}(T)z(T)dx-\int_{\Omega}z_{t}(0)z(0)dx\right]}_{J_{3}}-\underbrace{\int_{0}^{T}\int_{\Omega}\left[g(u^{1}_{t})-g(u^{2}_{t})\right]zdxds}_{J_{4}}.\end{aligned}$ (5.11)
From (5.3), we get
$2\int_{0}^{T}\|z_{t}\|^{2}ds\leq\frac{2}{\kappa_{2}}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds.$ (5.12)
From embedding $H^{1}\hookrightarrow H^{0}$ and using inequality (3.49), we have
$|-J_{3}|\leq\frac{1}{\lambda_{1}^{1/2}}\left[\,E_{z}(T)+E_{z}(0)\,\right].$ (5.13)
Still, using equality (LABEL:ASP-11) and estimates $|\mathrm{J}_{1}|$ and $|\mathrm{J}_{2}|$ (with $\tau=0$) and choosing $\delta=\frac{\lambda_{1}^{1/2}}{2}$, we obtain
$\begin{aligned} E_{z}(0)\leq&\;E_{z}(T)+\frac{3}{2}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds+\frac{K^{2}_{f}}{2\kappa_{2}}\int_{0}^{T}\|z\|^{2}ds\\ &+\frac{\lambda_{1}^{1/2}}{2}\int_{0}^{T}E_{z}(s)ds+C_{B}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds.\end{aligned}$ (5.14)
Then, substituting (5.14) in (5.13), we obtain
$\begin{aligned} |-J_{3}|\leq&\;\frac{2}{\lambda_{1}^{1/2}}E_{z}(T)+\frac{3}{2\lambda_{1}^{1/2}}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds+\frac{K^{2}_{f}}{2\kappa_{2}\lambda_{1}^{1/2}}\int_{0}^{T}\|z\|^{2}ds\\ &+\frac{1}{2}\int_{0}^{T}E_{z}(s)ds+C_{B}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds.\end{aligned}$ (5.15)
From Assumption (2.1) and (5.3), we have
$\begin{aligned} |-J_{4}|=&\;\left|-\int_{0}^{T}\int_{\Omega}\int_{0}^{1}g^{\prime}(\theta u^{1}_{t}+(1-\theta)u^{2}_{t})d\theta z_{t}zdxds\right|\\ \leq&\;\kappa_{1}\int_{0}^{T}\int_{\Omega}\int_{0}^{1}\left[\,1+|\theta u^{1}_{t}+(1-\theta)u^{2}_{t}|^{4}\,\right]d\theta|z_{t}||z|dxds\\ \leq&\;\kappa_{1}\int_{0}^{T}\|z_{t}\|\|z\|ds+8\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}||z|dxds\\ \leq&\;\int_{0}^{T}\|z_{t}\|^{2}ds+\frac{\kappa_{1}^{2}}{4}\int_{0}^{T}\|z\|^{2}ds+8\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}||z|dxds\\ \leq&\;\frac{1}{\kappa_{2}}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds+\frac{\kappa_{1}^{2}}{4}\int_{0}^{T}\|z\|^{2}ds+\underbrace{8\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}||z|dxds}_{J_{4}^{\prime}}\end{aligned}$ (5.16)
where from Hölder’s inequality with $\frac{1}{2}+\frac{1}{2}=1$ and $\frac{2}{3}+\frac{1}{3}=1$, embedding $H^{1}\hookrightarrow L^{6}(\Omega)$, (3.49), and (5.3), we have
$\begin{aligned} J^{\prime}_{4}\leq&\;8\kappa_{1}\int_{0}^{T}\left(\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}|^{2}dx\right)^{1/2}\left(\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z|^{2}dx\right)^{1/2}ds\\ \leq&\;4\kappa_{1}\int_{0}^{T}\left[\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}|^{2}dx+\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z|^{2}dx\right]\\ \leq&\;4\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}|^{2}dxds\\ &+4\kappa_{1}\int_{0}^{T}\left(\int_{\Omega}\left[\,|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\,\right]^{3/2}dx\right)^{2/3}\left(\int_{\Omega}|z|^{6}dx\right)^{1/3}ds\\ \leq&\;4\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}|^{2}dxds+4\kappa_{1}\int_{0}^{T}\left[\,|u^{1}_{t}\|^{4}_{6}+\|u^{2}_{t}\|^{4}_{6}\,\right]\|z\|^{2}_{6}ds\\ \leq&\;4\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}|^{2}dxds+\frac{4\kappa_{1}}{\lambda}\int_{0}^{T}\left[\,|u^{1}_{t}\|^{4}_{6}+\|u^{2}_{t}\|^{4}_{6}\,\right]\|\nabla z\|^{2}ds\\ \leq&\;4\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}|^{2}dxds+\frac{8\kappa_{1}}{\lambda}\int_{0}^{T}\left[\,|u^{1}_{t}\|^{4}_{6}+\|u^{2}_{t}\|^{4}_{6}\,\right]E_{z}(s)ds\\ \leq&\;4\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}|^{2}dxds+\frac{32\kappa_{1}^{2}}{\lambda^{2}}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds+\frac{1}{2}\int_{0}^{T}E_{z}(s)ds\\ \leq&\;\frac{40\kappa_{1}}{\kappa_{0}}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds+\frac{32\kappa_{1}^{2}}{\lambda^{2}}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds+\frac{1}{2}\int_{0}^{T}E_{z}(s)ds.\end{aligned}$ (5.17)
Replacing (5.12), (5.15), (5.16), and (5.17) in (5.11), we obtain that
$\begin{aligned} \int_{0}^{T}E_{z}(s)ds\leq&\;\frac{2}{\lambda_{1}^{1/2}}E_{z}(T)+C_{0}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds\\ &+C_{1}\int_{0}^{T}\|z\|^{2}ds+C^{\prime}_{B}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds,\end{aligned}$ (5.18)
where $C_{0}=\frac{2}{\kappa_{2}}+\frac{3}{2\lambda_{1}^{1/2}}+\frac{1}{\kappa_{2}}+\frac{40\kappa_{1}}{\kappa_{2}}$, $C_{1}=K_{f}+\frac{K_{f}^{2}}{2\kappa\lambda_{1}^{1/2}}+\frac{\kappa_{1}^{2}}{4}$, and $C^{\prime}_{B}=C_{B}+\frac{32\kappa_{1}^{2}}{\lambda^{2}}$. Combining (5.10) and (5.18) we obtain that
$\begin{aligned} \left(T-\frac{4}{\lambda_{1}^{1/2}}\right)E_{z}(t)+\left(1-\delta T\right)\int_{0}^{T}E(s)ds\leq\;2C_{0}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds\\ +\left(TC_{B,\delta}+2C^{\prime}_{B}\right)\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds+\left(2C_{1}+\frac{K_{f}T}{2\kappa_{2}}\right)\int_{0}^{T}\|z\|^{2}ds.\end{aligned}$ (5.19)
Using without loss of generality that $\frac{T}{2}-\frac{4}{\lambda_{1}^{1/2}}>0$ and taking $\delta$ small enough such that $1-\delta T\geq\frac{1}{2}$, it follows from (5.19) that
$\begin{aligned} TE_{z}(T)+\int_{0}^{T}E(s)ds\leq\;4C_{0}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds\\ +2\left(TC_{B,\delta}+2C^{\prime}_{B}\right)\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds+2\left(2C_{1}+\frac{K_{f}T}{2\kappa_{2}}\right)\int_{0}^{T}\|z\|^{2}ds.\end{aligned}$ (5.20)
From (5.9) with $\tau=0$, we have
$\begin{aligned} \frac{1}{2}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds\leq&\;E_{z}(0)-E_{z}(T)+\delta\int_{0}^{T}E_{z}(s)ds\\ &+C_{B,\delta}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds.\end{aligned}$ (5.21)
Substituting (5.21) into (5.20) and choosing $\delta=\frac{1}{2C_{0}}$, we obtain
$E_{z}(T)$ $\leq$ $\frac{2C_{0}}{T+2C_{0}}E_{z}(0)+\frac{4C_{0}\left(TC_{B,\delta}+2C^{\prime}_{B}\right)}{T+2C_{0}}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds$
$+\,\frac{4C_{0}\left(2C_{1}+\frac{K_{f}T}{2\kappa_{2}}\right)}{T+2C_{0}}\sup_{s\in[0,T]}\|z(s)\|^{2}.$
Reiterating the estimate on the intervals $[mT,(m+1)T]$ yields
$E_{z}((m+1)T)\leq\gamma E_{z}(mT)+C_{B}b_{m},\quad m=0,1,2,\cdots,$
with $0<\gamma\equiv\frac{2C_{0}}{T+2C_{0}}<1$ and $C_{B}=\frac{4C_{0}\left(TC_{B,\delta}+2C^{\prime}_{B}\right)}{T+2C_{0}}+\frac{4C_{0}\left(2C_{1}+\frac{K_{f}T}{2\kappa_{2}}\right)}{T+2C_{0}}$, where
$b_{m}\equiv\sup_{s\in[mT,(m+1)T]}\|z(s)\|^{2}+\int_{mT}^{(m+1)T}d(s;u^{1},u^{2})E_{z}(s)ds.$
This yields
$E_{z}(mT)\leq\gamma^{m}E_{z}(0)+c\sum_{l=1}^{m}\eta^{m-l}b_{l-1}.$
Since $\gamma<1$, using the same argument as in ([13], Remark 3.30) along with the definition of $b_{l}$ we obtain that there exists $\varrho>0$ such that
$E_{z}(t)\leq C_{1}e^{-\varrho t}E_{z}(0)+C_{2}\left[\,\sup_{0\leq s\leq t}\|z(s)\|^{2}+\int_{0}^{t}e^{-\varrho(t-s)}d(s,u^{1},u^{2})E_{z}(s)ds\,\right],$
for all $t\geq 0$. Therefore, applying Gronwall’s lemma we find
$E_{z}(t)\leq\left[\,C_{1}e^{-\varrho t}E_{z}(0)+C_{2}\sup_{0\leq s\leq t}\|z(s)\|^{2}\,\right]e^{C_{2}\int_{0}^{t}d(s,u^{1},u^{2})ds}.$
[…]
where
$b(t):=2C_{B}C_{1}e^{-\varrho t}e^{C_{2}\int_{0}^{t}d(s,u^{1},u^{2})ds}\quad\mbox{and}\quad c(t):=2C_{2}e^{C_{2}\int_{0}^{t}d(s,u^{1},u^{2})ds}.$ (5.22)
Thus, using that $d(s,u^{1},u^{2})=\|u^{1}_{t}\|^{6}_{6}+\|u^{2}_{t}\|^{6}_{6}\in L^{1}(0,t)$, uniformly in $t>0$ we obtain
$b(t)\in L^{1}(\mathbb{R}^{+})\quad\mbox{and}\quad\lim_{t\rightarrow+\infty}b(t)=0$
and $c(t)$ is bounded on $\mathbb{R}^{+}$ due to $L^{1}(\mathbb{R}^{+})$ integrability of $d(s,u^{1},u^{2})$. The proof of Proposition 5.1 is now complete. ∎
6 Proof of Theorem 2.6: Properties of Attractors
6.1 Proof of Part (I) and Part (III) in Theorem 2.6.
On the virtue of Theorem 2.5, we have established that under the condition $g^{\prime}(0)>0$ the system $(\mathcal{H},S_{t})$ is “quasi-stable”. Therefore, the finite dimension of the attractor $\mathfrak{A}$ (Part (I)) is guaranteed by direct application of Theorem 7.3. As to Part (II) we are in a position to obtain at this point only partial result as stated in (6.2). Indeed, it follows from Proposition 4.1 that the uniform estimate
[…]
is valid for any trajectory $\{U(t):t\in\mathbb{R}\}$ lying in the attractor $\mathfrak{A}$. In addition to the quasi-stability property of the system $(\mathcal{H},S_{t})$, using that the function $c(t)$ defined in (5.22) has the property $c_{\infty}=\sup_{t\in\mathbb{R}^{+}}c(t)<\infty$, it follows by direct application of Theorem 7.4 that any complete trajectory $\{U(t):t\in\mathbb{R}\}$ that belongs to the global attractor $\mathfrak{A}$ has the following regularity properties
$u_{t}\in L^{\infty}(\mathbb{R};H^{1})\cap C(\mathbb{R};H^{0}),\quad u_{tt}\in L^{\infty}(\mathbb{R};H^{0}).$
Moreover, there exists $R_{1}>0$ such that
[…]
where $R_{1}$ depends on the constant $c_{\infty}$. To complete the proof of Part (II), it remains to prove that $u\in L^{\infty}(\mathbb{R};H^{2})$ and $\|u(t)\|_{H^{2}}\leq R_{2}$ for all $t\in\mathbb{R}$ for some $R_{2}>0$. Since the proof of this result requires technical arguments, we carry out the proof separately in Section 6.2.∎
The proof of Part (III) is an application of Theorem 7.5. First, let $\mathfrak{B}$ be the absorbing set obtained in the Proposition 4.1. From Proposition 5.1, the system $(\mathcal{H},S_{t})$ is quasi-stable on $\mathfrak{B}$. Now, let us consider the spaces $\mathcal{H}_{-s}=H^{-s+1}\times H^{-s}$ with $0<s\leq 1$. Next, we consider the mapping $t\in[0,T]\mapsto S_{t}z\in\mathcal{H}_{-1}=H^{0}\times H^{-1}$. Let $U(t)=S_{t}z=z(t)$ be a weak solution with initial data $z\in\mathfrak{B}$. It follows from Theorem 2.2 that $U_{t}$ has regularity $U_{t}=(u_{t},u_{tt})\in L^{6/5}(0,T;\mathcal{H}_{-1})$. Then, we have
[…]
$\leq$ $C_{\mathfrak{B}}|t_{2}-t_{1}|^{1/6},\quad 0\leq t_{1}\leq t_{2}\leq T.$
This shows that for each $z\in\mathfrak{B}$, the map $t\mapsto S_{t}z$ is Hölder continuous in the extended space $\mathcal{H}_{-1}$ with exponent $\gamma=1/6$. Now, we assume that $0<s<1$. Then, using that $\mathcal{H}_{0}\hookrightarrow\mathcal{H}_{-s}\hookrightarrow\mathcal{H}_{-1}$, applying interpolation theorem in each component of $\mathcal{H}_{-s}$ and using the Hölder continuity in $\mathcal{H}_{-1}$, we obtain
[…]
$\leq$ $C_{\mathfrak{B}}|t_{2}-t_{1}|^{s/6},\quad\mbox{with}\quad s\in(0,1].$
This shows that $t\mapsto S_{t}z$ is Hölder continuous in the extended spaces $\mathcal{H}_{-s}$ for $s\in(0,1]$. Therefore, the existence of a generalized exponential attractor in $\mathcal{H}_{-s}$ with $0<s\leq 1$, follows from Theorem 7.5. ∎
6.2 Maximal Regularity of the Attractor-Completion of the Proof of Part II
To obtain this result we will use arguments similar to those used by Chueshov and Lasiecka in [12, Section 4.2], see aso [29]. The proof follows the following 3 steps. In the first, we prove the $H^{2}$ regularity of the stationary solutions of problem (LABEL:P); in the second, using the regularity of stationary solutions, we establish the smoothness of the trajectories for negative times $t\to$-$\infty$. Finally, in step 3, assuming the regularity obtained in step 2 for an initial data $U(T_{0})\in H^{2}$ with $T_{0}<0$, we propagate this regularity to all $t\in\mathbb{R}.$
Step 1: Regularity of stationary solutions
Lemma 6.1.
The stationary solutions $u$ of problem (LABEL:P) are bounded in $H^{2}$.
Proof.
Let $u$ be a stationary solution of (LABEL:P). Then from the definition of $\mathcal{N}$, we have
$-\Delta u+f(u)=h,\quad u\in H^{1},$ (6.3)
which by elliptic regularity, $u\in H^{2}$. Multiplying (6.3) by $u$ and integrating over $\Omega$, we get
$\|\nabla u\|^{2}+\int_{\Omega}f(u)udx=\int_{\Omega}hudx.$ (6.4)
From Assumption (2.5), we have
$\int_{\Omega}f(u)udx\geq-\frac{\nu}{2}\|u\|^{2}-C_{\nu}|\Omega|\geq-\frac{\nu}{\lambda_{1}}\|\nabla u\|^{2}-C_{\nu}|\Omega|.$ (6.5)
From Hölder inequality, embedding $H^{1}\hookrightarrow H^{0}$, and Young’s inequality, we have
$\left|\int_{\Omega}hudx\right|\leq\|h\|\|u\|\leq\|h\|\frac{1}{\lambda_{1}^{1/2}}\|\nabla u\|\leq\frac{1}{2\omega\lambda_{1}}\|h\|^{2}+\frac{\omega}{2}\|\nabla u\|^{2}.$ (6.6)
Using that $\omega=1-\frac{\nu}{\lambda_{1}}$ and combining (6.4), (6.5), and (6.6), we obtain that
$\|\nabla u\|^{2}\leq\frac{1}{\omega^{2}\lambda_{1}}\|h\|^{2}+\frac{2C_{\nu}|\Omega|}{\omega}.$ (6.7)
Now, multiplying (6.3) by $\Delta u$ and integrating over $\Omega$, we get
[…]
From Assumption (2.4) and embedding $H^{2}\hookrightarrow H^{1}$, we get
$-\int_{\Omega}f(u)\Delta udx=\int_{\Omega}f^{\prime}(u)|\nabla u|^{2}dx\geq-K_{f}\|\nabla u\|^{2}.$
where $K_{f}$ arises as in (3.25). From Hölder and Young inequalities, we have
$\left|-\int_{\Omega}h\Delta udx\right|\leq\frac{1}{2\omega}\|h\|^{2}+\frac{\omega}{2}\|\Delta u\|^{2}.$
Substituting the last two inequalities in (6.8) and using (6.7), we conclude that
$||\Delta u||^{2}\leq\frac{1}{\omega^{2}}\|h\|^{2}+\frac{2K_{f}}{\omega}\|\nabla u\|^{2}\leq\frac{1}{\omega^{2}}\|h\|^{2}+\frac{2K_{f}}{\omega}\left(\frac{1}{\omega^{2}\lambda_{1}}\|h\|^{2}+\frac{2C_{\nu}|\Omega|}{\omega}\right)=:\varrho.$ (6.9)
This completes the proof of the lemma. ∎
Step 2: Smoothness on negative time scale
Decomposition of the solutions.
Let $\{U(t)=(u(t),u_{t}(t)):t\in\mathbb{R}\}$ be a trajectory from the global attractor $\mathfrak{A}=\mathcal{M}(\mathcal{N})$ and let’s consider the following decomposition of the solution $u=w+z$ with $W=(w,w_{t})$ satisfying the problem
$\left\{\begin{array}[]{l}w_{tt}-\Delta w+g(w_{t})+f(w)=h,\quad\mbox{in}\quad\Omega\times(s,T)\\ w=0,\quad\mbox{on}\quad\Gamma,\\ (w(s),w_{t}(s))=(u^{*},0),\end{array}\right$. (6.13)
where $u^{*}\in H^{2}$ is a stationary solution of (LABEL:P), and $Z=(z,z_{t})$ satisfying the following problem
$\left\{\begin{array}[]{l}z_{tt}-\Delta z+g(u_{t})-g(w_{t})=-[\,f(u)-f(w)\,],\quad\mbox{in}\quad\Omega\times(s,T),\\ z=0,\quad\mbox{on}\quad\Gamma,\\ (z(s),z_{t}(s))=(u(s)-u^{*},u_{t}(s)).\end{array}\right$. (6.17)
Estimate 1.
The estimates performed below are “formal”. However, they can be easily justified by considering Galerkin approximations Taking the inner product of the Eq. in (6.13) with $\Delta w_{t}$ in $H^{0}$, yields
[…]
where $\mathcal{H}_{1}=H^{2}\times H^{1}$. On the other hand, taking the inner product of the Eq. in (6.13) with $\alpha\Delta w$, with $\alpha\in(0,1)$, we get
$\begin{aligned} &\alpha\frac{d}{dt}\int_{\Omega}\nabla w_{t}\nabla wdx+\alpha\|\Delta w(t)\|^{2}+\alpha\int_{\Omega}f^{\prime}(w)|\nabla w|^{2}dx+\alpha\int_{\Omega}h\Delta wdx\\ &=\alpha\|\nabla w_{t}(t)\|^{2}-\alpha\int_{\Omega}g^{\prime}(w_{t})\nabla w\nabla w_{t}dx.\end{aligned}$ (6.19)
Then, combining (6.18) and (6.19), we obtain
$\begin{aligned} &\frac{d}{dt}\mathcal{E}_{w}(t)+\alpha\mathcal{E}_{w}(t)+\frac{\alpha}{2}\|\Delta w(t)\|^{2}+\int_{\Omega}g^{\prime}(w_{t})|\nabla w_{t}|^{2}dx\\ &=\frac{3\alpha}{2}\|\nabla w_{t}(t)\|^{2}+\frac{1}{2}\int_{\Omega}f^{\prime\prime}(w)w_{t}|\nabla w|^{2}dx-\alpha\int_{\Omega}g^{\prime}(w_{t})\nabla w\nabla w_{t}dx+\alpha\int_{\Omega}\nabla w_{t}\nabla wdx.\end{aligned}$ (6.20)
where
[…]
Note that, from dissipative condition 2.3, Holder inequality, and embedding $H^{2}\hookrightarrow H^{1}$, we can estimate the functional $\mathcal{E}_{w}(t)$ from below as follows
$\begin{aligned} \mathcal{E}_{w}(t)\geq&\;\frac{1}{4}\|\Delta w(t)\|^{2}-\|h\|^{2}-\frac{K_{f}+\alpha^{2}}{2}\|\nabla w\|^{2},\end{aligned}$ (6.21)
where $K_{f}$ appears as in (3.25). So, we define the perturbed functional $\widetilde{\mathcal{E}}_{w}(t)$ by
$\widetilde{\mathcal{E}}_{w}(t):=\mathcal{E}_{w}(t)+\frac{K_{f}+\alpha^{2}}{2}\|\nabla w\|^{2}+\|h\|^{2},$
it follows from (6.21) that
$\widetilde{\mathcal{E}}_{w}(t)\geq\frac{1}{4}\|\Delta w(t)\|^{2}.$ (6.22)
Returning to (6.20), we obtain
$\begin{aligned} &\frac{d}{dt}\widetilde{\mathcal{E}}_{w}(t)+\alpha\widetilde{\mathcal{E}}_{w}(t)+\frac{\alpha}{2}\|\Delta w(t)\|^{2}+\int_{\Omega}g^{\prime}(w_{t})|\nabla w_{t}|^{2}dx\\ &=\alpha\|h\|^{2}+\frac{3\alpha}{2}\|\nabla w_{t}(t)\|^{2}+\alpha\int_{\Omega}\nabla w_{t}\nabla wdx+(K_{f}+\alpha^{2})\int_{\Omega}\nabla w_{t}\nabla wdx\\ &\quad+\frac{\alpha(K_{f}+\alpha^{2})}{2}\|\nabla w(t)\|^{2}+\frac{1}{2}\int_{\Omega}f^{\prime\prime}(w)w_{t}|\nabla w|^{2}dx-\alpha\int_{\Omega}g^{\prime}(w_{t})\nabla w\nabla w_{t}dx.\end{aligned}$ (6.23)
Now let’s estimate the terms on the right-hand side of (6.23). First, from embedding $H^{2}\hookrightarrow H^{1}$, (6.1), and (6.2), it follows:
$\alpha\|h\|^{2}+\frac{3\alpha}{2}\|\nabla w_{t}\|^{2}+\alpha\int_{\Omega}\nabla w_{t}\nabla wdx+(K_{f}+\alpha^{2})\int_{\Omega}\nabla w_{t}\nabla wdx+\frac{\alpha(K_{f}+\alpha^{2})}{2}\|\nabla w\|^{2}$
$\leq\alpha\|h\|^{2}+\frac{3\alpha R^{2}_{1}}{2}+\alpha R_{1}R+(K_{f}+\alpha^{2})R_{1}R+\frac{\alpha(K_{f}+\alpha)}{2}R^{2}=:C_{R,R_{1},h}.$
From Assumption (2.2), Holder inequality with $\frac{1}{2}+\frac{1}{6}+\frac{1}{3}=1$, embeddings $H^{2}\hookrightarrow H^{1}\hookrightarrow L^{6}(\Omega)$, (6.1), and (6.22), we have
$\frac{1}{2}\int_{\Omega}f^{\prime\prime}(w)w_{t}|\nabla w|^{2}dx$ $\leq$ $C\left[\,1+\|w(t)\|^{3}_{6}\,\right]\|w_{t}(t)\|_{6}\|\nabla w(t)\|^{2}_{6}$
$\leq$ $C\left[\,1+R^{3/2}\,\right]\|w_{t}(t)\|_{6}\|\Delta w(t)\|^{2}$
$\leq$ $\frac{\alpha}{2}\widetilde{\mathcal{E}}_{w}(t)+\mathbf{d}(t,w)\widetilde{\mathcal{E}}_{w}(t),$
where $\mathbf{d}(t,w):=C_{R}\|w_{t}(t)\|^{6}_{6}$. Finally, from Young’s inequality, Assumption (2.1), Holder inequality with $\frac{2}{3}+\frac{1}{3}=1$, embedding $H^{1}\hookrightarrow L^{6}(\Omega)$, and (6.2), we have
$-\alpha\int_{\Omega}g^{\prime}(w_{t})\nabla w\nabla w_{t}dx$ $\leq$ $\int_{\Omega}g^{\prime}(w_{t})|\nabla w_{t}|^{2}dx+\frac{\alpha^{2}}{4}\int_{\Omega}g^{\prime}(w_{t})|\nabla w|^{2}dx$
$\leq$ $\int_{\Omega}g^{\prime}(w_{t})|\nabla w_{t}|^{2}dx+\frac{\alpha^{2}\kappa_{1}}{4}\int_{\Omega}\left[1+|w_{t}|^{4}\right]|\nabla w|^{2}dx$
$\leq$ $\int_{\Omega}g^{\prime}(w_{t})|\nabla w_{t}|^{2}dx+\alpha^{2}C^{\prime}_{R_{1}}\|\Delta w(t)\|^{2}.$
Thus, using the last three inequalities, we conclude from (6.23) that
$\frac{d}{dt}\widetilde{\mathcal{E}}_{w}(t)+\frac{\alpha}{2}\widetilde{\mathcal{E}}_{w}(t)+\alpha\left(\frac{1}{2}-\alpha C^{\prime}_{R_{1}}\right)\|\Delta w(t)\|^{2}\leq\mathbf{d}(t,w)\widetilde{\mathcal{E}}_{w}(t)+C_{R,R_{1},h},$ (6.24)
Now, choosing $\alpha$ small enough such that $\frac{1}{2}-\alpha C^{\prime}_{R_{1}}\geq 0$, it follows from (6.24) that
$\begin{aligned} \frac{d}{dt}\widetilde{\mathcal{E}}_{w}(t)+\frac{\alpha}{2}\widetilde{\mathcal{E}}_{w}(t)\leq\mathbf{d}(t,w)\widetilde{\mathcal{E}}_{w}(t)+C_{R,R_{1},h}.\end{aligned}$ (6.25)
Multiplying (6.25) by the integrating factor $e^{\frac{\alpha}{2}t}$ and integrating from $s$ to $t$, we obtain
$\begin{aligned} \widetilde{\mathcal{E}}_{w}(t)e^{\frac{\alpha}{2}t}\leq e^{\frac{\alpha}{2}s}\widetilde{\mathcal{E}}_{w}(s)+\int_{s}^{t}\mathbf{d}(\tau,w)e^{\frac{\alpha}{2}\tau}\widetilde{\mathcal{E}}_{w}(\tau)d\tau+\frac{2C_{R_{1},R_{2},h}}{\alpha}e^{\frac{\alpha}{2}t}.\end{aligned}$ (6.26)
Using that $\mathbf{d}(\cdot,w)\in L^{1}(s,t)$ and (6.22), it follows from (6.26) that
$\begin{aligned} \|\Delta w(t)\|^{2}e^{\frac{\alpha}{2}t}\leq 4e^{\|\mathbf{d}\|_{L^{1}}}\left[\,e^{\frac{\alpha}{2}s}\widetilde{\mathcal{E}}_{w}(s)+\frac{2C_{R,R_{1},h}}{\alpha}e^{\frac{\alpha}{2}t}\right].\end{aligned}$ (6.27)
Note that, from (6.9) (Lemma 6.1) [regularity of steady states], we have
$\widetilde{\mathcal{E}}_{w}(s)\leq C\|\Delta u^{*}\|^{2}+\frac{1}{2}\|h\|^{2}\leq C\varrho+\frac{1}{2}\|h\|^{2}=:\varrho_{0}.$ (6.28)
Thus, taking the limit in (6.27) with $s\to$-$\infty$, we have
$\begin{aligned} \|\Delta w(t)\|^{2}e^{\frac{\alpha}{2}t}\leq\frac{8C_{R,R_{1},h}e^{\|\mathbf{d}\|_{L^{1}(s,t)}}}{\alpha}e^{\frac{\alpha}{2}t}.\end{aligned}$
This implies that
$\begin{aligned} \|\Delta w(t)\|^{2}\leq\frac{8C_{R,R_{1},h}e^{\|\mathbf{d}\|_{L^{1}(s,t)}}}{\alpha}.\end{aligned}$ (6.29)
As for (6.28), $\widetilde{\mathcal{E}}_{w}(s)\leq\varrho_{0}$ independent on $s$, we conclude from (6.29) that the full trajectory $\{W(t)=(w(t),w_{t}(t)):t\in\mathbb{R}\}$ has $H^{2}\times H^{1}$ regularity.
Estimate 2.
Our goal now is to show that the original trajectory $U(t)$ coincides with $W(t)$ for $t\in(-\infty,T_{0}]$ for some $T_{0}<0$. In fact, multiplying (6.17) by $z_{t}$ and integrating over $\Omega\times[s,t]$, we have
$\begin{aligned} &E_{z}(t)+\int_{s}^{t}\int_{\Omega}\left[\,g(u_{t})-g(w_{t})\,\right]z_{t}dxd\tau=E_{z}(s)\\ &\quad+\underbrace{K_{f}\int_{s}^{t}\int_{\Omega}z_{t}zdxd\tau}_{\mathrm{J}_{1}}+\underbrace{\frac{1}{2}\int_{s}^{t}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\theta u+(1-\theta)w)[u_{t}+(1-\theta)w_{t}]d\theta|z|^{2}dxd\tau}_{\mathrm{J}_{2}}.\end{aligned}$ (6.30)
where
$E_{z}(t)=\frac{1}{2}\|z_{t}\|^{2}+\frac{1}{2}\|\nabla z\|^{2}+\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime}(\theta u+(1-\theta)w)d\theta|z|^{2}dx+\frac{K_{f}}{2}\|z\|^{2}.$
From embedding $H^{1}\hookrightarrow H^{0}$, we have
$\left|\mathrm{J}_{1}\right|\leq\frac{K_{f}}{\lambda_{1}^{1/2}}\int_{s}^{T}E_{z}(\tau)d\tau.$
From Hölder inequality with $\frac{2}{3}+\frac{1}{6}+\frac{1}{6}=1$, embedding $H^{1}\hookrightarrow L^{6}(\Omega)$, and uniform estimates (6.2), we have
$\left|\mathrm{J}_{2}\right|\leq C_{R,R_{1}}\int_{s}^{t}E_{z}(\tau)d\tau.$
Substituting $|\mathrm{J}_{1}|$ and $|\mathrm{J}_{2}|$ in (6.30), we get
$E_{z}(t)\leq E_{z}(s)+C^{\prime}_{R,R_{1}}\int_{s}^{t}E_{z}(\tau)d\tau.$ (6.31)
Applying Gronwall’s lemma in (6.31), we obtain
$E_{z}(t)\leq e^{C^{\prime}_{R_{1},R_{2}}(t-s)}E_{z}(s).$ (6.32)
Taking $t=s+r$, with $r>0$, from (6.32), we have
$E_{z}(s+r)\leq e^{C^{\prime}_{R,R_{1}}r}E_{z}(s).$ (6.33)
[…]
From (6.34) and (6.17) , we conclude that
[…]
where we have used $Z(s)\rightarrow 0,s\rightarrow$-$\infty$. That is, for any $\varepsilon>0$ there exists a $T_{0}=s_{0}+r<0$ such that
[…]
This proves that the trajectories $U(t)$ and $W(t)$ coincide for $t\in(-\infty,T_{0}]$.
Step 3: Forward propagation of the regularity
Let us now consider the problem (LABEL:P) with initial data $U(T_{0})\in H^{2}\times H^{1}$ for $T_{0}<0$ as obtained in Step 2. That is,
$\left\{\begin{array}[]{l}u_{tt}-\Delta u+g(u_{t})+f(u)=h,\quad\mbox{in}\quad\Omega\times(T_{0},\infty)\\ w=0,\quad\mbox{on}\quad\Gamma,\\ (u(T_{0}),u_{t}(T_{0}))=(u^{*}_{0},u^{*}_{1})\in H^{2}\times H^{1}.\end{array}\right$.
Proceeding analogously to Estimate 1 of Step 2, we find the following inequality (analogous to (6.27))
$\|\Delta u(t)\|^{2}\leq 4e^{\|d\|_{L^{1}(T_{0},t)}}e^{-\varpi(t-T_{0})}\widetilde{\mathcal{E}}_{u}(T_{0})+4e^{\|d\|_{L^{1}(T_{0},t)}}K_{R,R_{1},h},\quad\forall t\geq T_{0},$
[…]
$\|u(t)\|_{H^{2}}\leq R_{2},\quad\forall t\in\mathbb{R}.$
This concludes the proof of maximal regularity, hence of Part (II).
7 Appendix: Definitions and Some Relevant Results Pertaining to the Theory of Abstract Dynamical systems
The results of the abstract theory of attractors presented below can be found in [13, 14].
7.1 Basic Definitions
Let $(X,S_{t})$ be a dynamical system given by a $C_{0}$-semigroup $S_{t}$ on a complete metric space $X$.
Definition 7.1.
A bounded set $\mathfrak{B}\subset X$ is an absorbing set for $S_{t}$ if for any bounded set $\mathcal{B}\subset X$, there exists $t_{\mathcal{B}}\geq 0$ such that
$S_{t}\mathfrak{B}\subset\mathfrak{B},\quad\forall\,t\geq t_{\mathcal{B}},$
which defines $(X,S_{t})$ as a dissipative dynamical system.
Definition 7.2.
$(X,S_{t})$ is said to be asymptotically smooth in $X$ if for any bounded positive invariant set $\mathcal{B}\subset X$, there exists a compact set $K\subset\overline{\mathcal{B}}$, such that
$\mbox{dist}(S_{t}\mathcal{B},K)=0\quad\mbox{as}\quad t\to\infty.$
Definition 7.3.
A bounded closed set $\mathfrak{A}\subset X$ is said to be a global attractor of $(X,S_{t})$ iff the following properties hold: (i) $\mathfrak{A}\subset X$ is an invariant set, that is, $S_{t}\mathfrak{A}=\mathfrak{A}$ for all $t\geq 0$; (ii) $\mathfrak{A}\subset X$ is uniformly attracting, that is, for all bounded $D\subset X$
$\mbox{dist}(S_{t}D,\mathfrak{A})=\sup_{x\in S_{t}D}\inf_{y\in\mathfrak{A}}d(x,y)\to 0\quad\mbox{as}\quad t\to\infty.$
Definition 7.4.
Let $\mathcal{N}$ be the set of stationary points of $(X,S_{t})$:
$\mathcal{N}=\{v\in X:S_{t}v=v\;\;\mbox{for all}\;\;t\geq 0\}.$
We define the unstable manifold $\mathrm{M}^{u}(\mathcal{N})$ emanating from set $\mathcal{N}$ as a set of all $y\in X$ such that there exists a full trajectory $\Upsilon=\{u(t):t\in\mathbb{R}\}$ with the properties
$u(0)=y\quad\mbox{and}\quad\lim_{t\rightarrow-\infty}\mbox{dist}_{X}(u(t),\mathcal{N})=0.$
Definition 7.5.
$(X,S_{t})$ is said to be gradient if it possesses a strict Lyapunov function, i.e. there exists a continuous functional $\Phi(y)$ defined on $X$ such that
  • (i)
    the function $t\rightarrow\Phi(S_{t}y)$ is nonincreasing for any $y\in X$, and
  • (ii)
    the equation $\Phi(S_{t}y)=\Phi(y)$ for all $t>0$ and for some $y\in X$ implies that $S_{t}y=y$ for all $t>0$, i.e. $y$ is a stationary point of $(X,S_{t})$.
Definition 7.6.
The fractal dimension of a compact set $K\subset X$ is defined by
$\mbox{dim}K_{f}^{X}=\lim\sup_{\varepsilon\rightarrow 0}\frac{\ln(n(K,\varepsilon))}{\ln(1/\varepsilon)},$
where $n(X,\varepsilon)$ is the minimal number of closed balls of the radius $\varepsilon$ which cover the set $K$.
Definition 7.7.
A compact set $\mathfrak{A}_{\exp}\subset X$ is said to be inertial (or a fractal exponential attractor) of the dynamical system $(X,S_{t})$ iff $\mathfrak{A}_{\exp}$ is a positively invariant set of finite dimension and for every bounded set $D\subset X$ there exist positive constants $t_{D},C_{D}$ and $\gamma_{D}$ such that
$d_{X}\{S_{t}D|\mathfrak{A}_{\exp}\}\equiv\sup_{x\in D}\mbox{dist}_{X}\left(S_{t}x,\mathfrak{A}_{\exp}\right)\leq C_{D}e^{-\gamma_{D}(t-t_{D})},\quad t\geq t_{D}.$
Definition 7.8.
Let $X,Y$ be two reflexive Banach spaces with $X$ compactly embedded in $Y$ and set $H=X\times Y$. Consider the dynamical system $(H,S_{t})$ given by an evolution operator
$S_{t}z=(u(t),u_{t}(t)),\qquad z=(u_{0},u_{1})\in H,$
where the functions $u$ and $\theta$ possess the regularity
$u\in C(\mathbb{R}^{+};X)\cap C^{1}(\mathbb{R}^{+};Y),\quad$
Then one says that $(H,S_{t})$ is quasi-stable on a set $\mathcal{B}\subset H$ if there exist a compact seminorm $n_{X}$ on $X$ and nonnegative scalar functions $a(t)$ and $c(t)$ locally bounded in $0,\infty)$, and $b(t)\in L^{1}(\mathbb{R}^{+})$ with $\lim_{t\to\infty}b(t)=0,$ such that
$\|S(t)z^{1}-S(t)z^{2}\|_{H}^{2}$ $\leq$ $a(t)\|z^{1}-z^{2}\|_{H}^{2},$ (7.1)
$\|S(t)z^{1}-S(t)z^{2}\|_{H}^{2}$ $\leq$ $b(t)\|z^{1}-z^{2}\|_{H}^{2}+c(t)\sup_{0<s<t}\left[n_{X}(u^{1}(s)-u^{2}(s))\right]^{2},$ (7.2)
for any $z^{1},z^{2}\in\mathcal{B}$. The inequality (7.2) is often called quasi-stability inequality.
7.2 Abstract Results
We now present the statements of the abstract results used in the proof of Theorem 2.4. In order, the next theorem guarantees that a dissipative system has a compact global attractor if and only if it is asymptotically compact.
Theorem 7.1.
[14, Theorem 7.2.3] Any dissipative asymptotically smooth dynamical system $(X,S_{t})$ in a Banach space $X$ possesses a unique compact global attractor $\mathfrak{A}$. This attractor is connected set and can be described as set of all bounded full trajectories.
The following result gives the characterization of the attractor for gradient systems.
Theorem 7.2.
[14, Theorem 7.5.6] Let a dynamical system $(X,S_{t})$ possess a compact global attractor $\mathfrak{A}$. Assume that there exists a strict Lyapunov function on $\mathfrak{A}$. Then $\mathfrak{A}=\mathrm{M}^{u}(\mathcal{N})$. Moreover, the global attractor $\mathfrak{A}$ consists of full trajectories $\Upsilon=\{u(t):t\in\mathbb{R}\}$ with the properties
$\lim_{t\rightarrow-\infty}\mbox{dist}_{X}(u(t),\mathcal{N})=0\quad\mbox{and}\quad\lim_{t\rightarrow+\infty}\mbox{dist}_{X}(u(t),\mathcal{N})=0.$
The next three results show that quasi-stable systems enjoy nice properties that include both finite-dimensional, smoothness, and existence of exponential attractor.
Theorem 7.3.
[14, Theorem 7.9.6] Assume that the dynamical system $(H,S_{t})$ possess a compact global attractor $\mathfrak{A}$ and is quasi-stable on $\mathfrak{A}$. Then the attractor $\mathfrak{A}$ of has a finite fractal dimension $dim^{H}_{f}\mathfrak{A}$.
Theorem 7.4.
[14, Theorem 7.9.8] Assume that the dynamical system $(H,S_{t})$ possess a compact global attractor $\mathfrak{A}$ and is quasi-stable on $\mathfrak{A}$. Moreover, we assume that (7.2) holds with the function $c(t)$ possessing the property $c_{\infty}=\sup_{t\in\mathbb{R}^{+}}c(t)<\infty$. Then any full trajectory $\{(u(t);u_{t}(t);\theta(t)):t\in\mathbb{R}\}$ that belongs to the global attractor enjoys the following regularity properties,
$u_{t}\in L^{\infty}(\mathbb{R};X)\cap C(\mathbb{R};Y),\quad u_{tt}\in L^{\infty}(\mathbb{R};Y),\quad\theta\in L^{\infty}(\mathbb{R};Z).$
Moreover, there exists $R>0$ such that
$|u_{t}(t)|_{X}^{2}+|u_{tt}(t)|_{Y}^{2}+|\theta_{t}(t)|_{Z}^{2}\leq R^{2},\quad t\in\mathbb{R},$
where $R$ depends on the constant $c_{\infty}$, on the semigroup $\eta_{X}$ in Definition [14, Definition 7.9.2], also on the embedding properties of $X$ into $Y$.
Theorem 7.5.
[14, Theorem 7.9.9] Assume that the dynamical system $(H,S_{t})$ is dissipative and quasi-stable on some bounded absorbing set $\mathfrak{B}$. We also assume that there exists a space $\widetilde{H}\subset H$ such that $t\mapsto S_{t}z$ is Hölder continuous in $\widetilde{H}$ for every $z\in\mathfrak{B}$, that is, there exist $0<\gamma\leq 1$ and $C_{\mathfrak{B},T}>0$ such that
$|S_{t_{1}}z-S_{t_{2}}z|_{\widetilde{H}}\leq C_{\mathfrak{B},T}|t_{1}-t_{2}|^{\gamma},\quad t_{1},t_{2}\in[0,T],\;z\in\mathfrak{B}.$
Then, the dynamical system $(H,S_{t})$ possesses a (generalized) fractal exponential exponential attractor whose dimension is finite in the space $\widetilde{H}$.